### v6

# Towards Federated Graph Learning in One-shot Communication

Abstract
Federated Graph Learning (FGL) has emerged as a promising paradigm for breaking data silos among distributed private graphs. In practical scenarios involving heterogeneous distributed graph data, personalized Federated Graph Learning (pFGL) aims to enhance model utility by training personalized models tailored to client needs. However, existing pFGL methods often require numerous communication rounds under heterogeneous graphs, leading to significant communication overhead and security concerns. While One-shot Federated Learning (OFL) enables collaboration in a single round, existing OFL methods are designed for image-centric tasks and are ineffective for graph data, leaving a critical gap in the field. Additionally, personalized models derived from existing methods suffer from bias, failing to effectively generalize to the minority. To address these challenges, we propose the first One-shot personalized Federated Graph Learning method (O-pFGL) for node classification, compatible with Secure Aggregation protocols for privacy preservation. Specifically, for effective graph learning in one communication round, our method estimates and aggregates class-wise feature distribution statistics to construct a global pseudo-graph on the server, facilitating the training of a global graph model. To mitigate bias, we introduce a two-stage personalized training approach that adaptively balances local personal information and global insights from the pseudo-graph, improving both personalization and generalization. Extensive experiments on 12 multi-scale graph datasets demonstrate that our method significantly outperforms state-of-the-art baselines across various settings.
1 Introduction
Graphs are widely employed to model complex relationships between entities across a variety of domains [20, 25, 3]. While various algorithms and models [16, 26, 50] have been proposed, most of them assume that graph data from different sources are managed centrally. However, collecting and managing such graph data is often costly, impractical, and poses significant privacy risks in many sensitive scenarios [14].
Figure 1: (a) A toy example explains the major classes and minor classes. (b) The empirical analysis of a client on the CiteSeer dataset under the Louvain partition to illustrate L3. Left: Existing methods result in a biased model that neglects the minority. Middle: The performance of existing methods and our method in class 1 with 50 samples. Right: The performance of existing methods and our method in class 2 with 14 samples. Our method effectively mitigates the performance gap.
To address the tension between the need for vast datasets and the growing demand for privacy protection, Federated Graph Learning (FGL) has emerged as a promising solution, which enables collaborative training on distributed graph data without requiring centralized management [37, 17]. While FGL shows its potential, it faces two major limitations: (1) Typical FGL struggles to effectively personalize models for joint improvement across heterogeneous clients’ graph data [2]. This heterogeneity arises from variations of both node and structural properties, with adjacent nodes influencing each other, further complicating personalization. (2) FGL requires iterative communication between clients and the server, raising efficiency concerns due to communication overhead [28, 8] and security concerns (e.g. man-in-the-middle attacks [47, 36, 67]). These limitations can be mitigated by integrating personalized federated graph learning (pFGL) and one-shot federated learning (OFL), which offer novel solutions to improve communication efficiency and model personalization.
To address these challenges, we explore the integration of pFGL and OFL. The pFGL methods allow for personalized model training by focusing on individual client graph data. OFL methods limit communication to a single round, thereby reducing the communication overhead and circumventing security risks such as man-in-the-middle attacks [18, 62]. However, both existing pFGL and OFL methods face challenges in FGL. Specifically, they exhibit the following limitations:
L1: Ineffectiveness for graph learning in one communication round. Existing pFGL methods [2, 33, 69] rely on iterative communication for optimization, making them ineffective within one communication round. Most methods also fail to support heterogeneous models, as they primarily rely on aggregating model parameters. Meanwhile, existing OFL methods are incompatible with graph data since they focus primarily on image-centric tasks. The ensemble-based OFL methods [62, 9] can generate images from ensemble models, but struggle with graph data due to the inter-dependencies between nodes. The distillation-based OFL faces the problems of distilling inter-related graphs. The generative-based OFL methods [21, 55, 54] rely on pre-trained generative models (e.g., Stable Diffusion [38]), which are impractical for graph generation.
L2: Incompatibility with Secure Aggregation. Secure Aggregation is a widely used technique [4, 41] designed to safeguard client privacy during the federated aggregation process. However, most pFGL and OFL methods require clients to upload raw model parameters or auxiliary information independently, which complicates the aggregation process (e.g., ensemble or similarity calculation) and hinders the implementation of privacy-preserving weighted averages.
L3: Imbalance between personalization and local generalization. Local graph data often exhibit a significant imbalance in both quantity and topology, as illustrated in Figure 1. Personalized models tend to be biased toward major classes (with larger quantities and high homophily to form compact communities), while neglecting minor classes with fewer samples and often surrounded by nodes from major classes. This imbalance leads to a degradation in model performance, particularly in underrepresented classes. A detailed empirical analysis is provided in Figure 1, which shows that existing personalized methods result in heavy bias toward majority classes (class 3 and class 5, with a total of 180 samples) while failing to adequately represent minority classes (class 1 and class 2, with only 64 samples), highlighting the limitations of current methods in handling imbalanced graph data.
To overcome these limitations, we propose a One-shot personalized Federated Graph Learning method (O-pFGL). Our method supports model heterogeneity and Secure Aggregation protocols for privacy preservation. Specifically, for L1 and L2, clients upload estimated feature distribution statistics, augmented by our proposed homophily-guided reliable node expansion strategy, instead of raw model parameters. The server securely aggregates these statistics to generate a global pseudo-graph, which is subsequently utilized to train personalized models on clients, eliminating the need for direct model parameter exchanges and preventing model theft. To address L3, we propose a two-stage personalized training approach. In the first stage, clients train a global generalized model based on the pseudo-graph, which serves as both a fixed teacher capturing global knowledge and a start-point for subsequent fine-tuning. In the second stage, clients fine-tune the global model on local graph data with the node-adaptive distillation, which leverages the global knowledge from the teacher model to counteract biases toward the majority, ensuring effective learning for the minority and achieving a robust balance between personalization and generalization.
Our method significantly outperforms state-of-the-art baselines through comprehensive experiments across various settings, including different graph scales, partition strategies, learning paradigms, and model architectures. Experimental results demonstrate the superiority of our method in achieving a one-shot personalized federated graph learning solution that balances efficiency, privacy, and personalization. To sum up, our contributions are listed as follows:
Bridge the Gap. To the best of our knowledge, we are the first to investigate the one-shot personalized federated graph learning problem under both graph data heterogeneity and model heterogeneity.
Novel Solution. We propose a novel method that aggregates client statistics rather than model parameters to generate a global pseudo-graph, which is used for two-stage personalized training with node-adaptive distillation. Our method is compatible with Secure Aggregation and protects models’ intellectual property.
SOTA Performance. We conduct extensive experiments on 12 multi-scale graph datasets across various settings and demonstrate the superiority of our method.
2 Related Works
2.1 Federated Graph Learning
From the graph level [52, 43], each client possesses multiple independent graphs (e.g., molecular graphs) and targets at graph-level tasks. From the subgraph level, each client possesses a subgraph of an implicit global graph. FedPUB [2] and FedGTA [33] personalize models by similarity of functional embedding and mixed moments of neighbor features. AdaFGL [32] studies the structure non-IID problem. FedTAD [69] performs data-free distillation on the server. To complete missing connections between subgraphs, FedSage [63] and FedDEP [64] additionally train a neighborhood generator. To better represent the minority in local graph data, FedSpray [15] learns class-wise structure proxies to mitigate biased neighborhoods. But it needs numerous communication rounds for optimization. FedLoG [30] synthesizes global data from clients’ synthetic nodes and generalizes the local training. Detailed description and critical analysis are in the Appendix A.6.
2.2 One-shot Federated Learning
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into: (1) Ensemble-based methods [18, 62, 9], (2) Distillation-based methods [66, 42], (3) Generative-based methods [21, 54, 55]. However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning. Detailed description and critical analysis are in the Appendix A.6.
Figure 2: The overall pipeline of our proposed method: O-pFGL.
3 Preliminaries
3.1 Notations and Definitions
Consider a federated graph learning setting with $m$ clients, where $k$-th client possesses the local undirected graph $G_{k}(\mathcal{V}_{k},\mathcal{E}_{k})$ with $|\mathcal{E}_{k}|$ edges and $|\mathcal{V}_{k}|=N_{k}$ nodes. Node features matrix is $\mathbf{X}_{k}=\{\mathbf{x}_{1},...,\mathbf{x}_{n_{k}}\}\in\mathbb{R}^{|\mathcal{V}_{k}|\times f}$, where $f$ is the feature dimension and $\mathbf{x}_{i}$ is the feature of node $v_{i}$. The adjacency matrix is $\mathbf{A}_{k}$ and the augmented normalized adjacency matrix is $\tilde{\mathbf{A}}_{k}$. The labels of nodes in its local graph are denoted by $\mathbf{Y}_{k}=\{y_{1},...,y_{n_{k}}\}\in\mathbb{R}^{|\mathcal{V}_{k}|\times C}$, where $C$ is the number of distinct classes. In the semi-supervised node classification task [31], nodes on the $k$-th client are partitioned into labeled nodes $\mathcal{V}_{k,L}$ (i.e., nodes in the training set) and unlabeled nodes $\mathcal{V}_{k,U}$ (i.e., their labels are unknown during training). The corresponding labels are $\mathbf{Y}_{k,L}$ and $\mathbf{Y}_{k,U}$, respectively. The neighborhood of node $v_{i}$ in $G_{k}(\mathcal{V}_{k},\mathcal{E}_{k})$ is defined as $\mathcal{N}_{v_{i}}=\{u|\exists e_{u,v_{i}}\in\mathcal{E}_{k}\}$, where edge $e_{u,v_{i}}$ connects node $u$ and $v_{i}$.
Definition 3.1.
Node Degree. For node $v_{i}$ in an undirected graph $G(\mathcal{V},\mathcal{E})$, the node degree $d(v_{i})$ is the number of edges connected to $v_{i}$. Formally, $d(v_{i})=|\{e_{v_{i},u}|,u\in\mathcal{V}\}|$.
Definition 3.2.
Node Homophily. For a labeled node $v_{i}\in\mathcal{V}_{L}$ in an undirected graph $G(\mathcal{V},\mathcal{E})$, the node homophily $h_{node}(v_{i})$ quantifies the proportion of its labeled neighbors that share the same label:
$h_{node}(v_{i})=\frac{\left|\{u|(u\in\mathcal{N}_{v_{i}}\cap\mathcal{V}_{L})\text{and}(y_{v_{i}}=y_{u})\}\right|}{|\mathcal{N}_{v_{i}}\cap\mathcal{V}_{L}|}.$ (1)
Definition 3.3.
Accumulated Class Homophily. The accumulated class homophily for class $c$ is defined as:
$H(c)=\sum_{u\in\mathcal{N}_{c}}h_{node}(u),$ (2)
where $\mathcal{N}_{c}=\{v_{i}\in\mathcal{V}_{L}|y_{i}=c\}$ is the set of labeled nodes belonging to class $c$. Note that $H(c)$ considers both the quantity and topology of nodes. A higher $H(c)$ indicates nodes of class $c$ are numerous and exhibit strong homophily, making them more predictable for a vanilla graph model [31].
Definition 3.4.
Class-aware Distillation Factor.
$w_{dist}(c)=\frac{1}{1+\log(H(c)+1)},$ (3)
which can be organized into a vector $\mathbf{w}_{dist}\in\mathbb{R}^{C}$ where the $i$-th entry corresponds to $w_{dist}(i)$ for class $i$. $\mathbf{w}_{dist}$ is used to determine node-adaptive distillation factor $\gamma_{i}$, as described in Section 4.4.
3.2 Problem Formulation
Assume a situation where a whole graph is distributed on $m$ clients, each of which possesses its local non-overlapped subgraph. We formulate the one-shot personalized federated graph learning for the node classification problem as:
$\min\sum_{k=1}^{m}L(M_{k}(\mathbf{A}_{k},\mathbf{X}_{k}),\mathbf{Y}_{k,U}),$ (4)
This objective aims to minimize the aggregated generalization loss $L$ on their unlabeled node set with personalized models $M_{k}$ across all clients within a single upload-download communication round.
4 Proposed Method
4.1 Overall Pipeline
The overall pipeline of our method is illustrated in Figure 2, which comprises 5 steps. (1) Each client estimates the class-wise feature distribution and computes corresponding statistics on its local graph data; (2) Each client uploads its locally estimated class-wise statistics to the server; (3) The server aggregates the uploaded statistics by weighted average to recover the global class-wise feature distribution. Based on the recovered distribution, the server generates a small-size global pseudo-graph; (4) The server distributes the generated global pseudo-graph to clients; (5) Each client performs two-stage personalized training using both local graph data and the downloaded global pseudo-graph, deriving a model with improved personalization and generalization.
Specifically, in the first step, we propose a homophily-guided reliable node expansion strategy to augment the distribution estimation. In the third step, we adapt techniques from graph condensation to facilitate the global pseudo-graph generation. In the fifth step, we address the critical bias issue in personalization by proposing a two-stage personalized training process with node-adaptive distillation. The details are explained in the following sections.
4.2 Feature Distribution Estimation
To estimate the class-wise feature distribution locally, each client propagates node features on local graph data. Specifically, on client $k$, the propagated features are:
$\mathbf{X}^{prop}_{k}=\|_{i=0}^{h}\tilde{\mathbf{A}}_{k}^{i}\mathbf{X}_{k},$ (5)
[…]
$\boldsymbol{\mu}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}|}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}}\mathbf{x}^{prop}_{i},\quad\boldsymbol{s^{2}}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}|-1}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}}(\mathbf{x}^{prop}_{i}-\boldsymbol{\mu}_{k}^{c})^{2},$ (6)
where $\mathcal{V}_{k,L}^{c}$ denotes the labeled nodes of class $c$ on client $k$. The statistics are estimated only for classes with sufficient labeled nodes, determined by a client-specified threshold.
However, the estimation may lack accuracy due to the limited number of labeled nodes. To address this, we propose a Homophily-guided Reliable node Expansion (HRE) strategy as a plugin to further augment the estimation process on homophilic graphs. In HRE, we first apply Label Propagation [26] to obtain the soft labels $\tilde{\mathbf{y}}\in\mathbb{R}^{C}$ for unlabeled nodes. We denote $c_{i}^{\prime}=\arg\max_{0\leq i<C}\tilde{\mathbf{y}}(i)$, then identify reliable soft labels based on the following criteria: (C1) High Node Degree: $d(v_{i})\geq d_{th}$; (C2) High Confidence: $\tilde{\mathbf{y}}_{i}(c_{i}^{\prime})\geq f_{th}$; (C3) High Class Homophily: $c_{i}^{\prime}\in topK(H)$, indicating the node is likely to share the same label with its neighbors. A node with more neighbors could reduce the variance in predictions. With these criteria, the reliable nodes $\mathcal{V}_{k,r}$ set with inferred true labels is:
$\mathcal{V}_{k,r}=\{v_{i}|(v_{i}\in\mathcal{V}_{k,U})\wedge(d(v_{i})\geq d_{th})\wedge(\tilde{\mathbf{y}}_{i}(c^{\prime})\geq f_{th})\wedge(c^{\prime}\in topK(H))\},$ (7)
where $topK(H)$ is the set of classes corresponding to the top $K$ largest values of accumulated class homophily $H(c)$. $K,f_{th}$ and $d_{th}$ are client-defined hyper-parameters. The reliable nodes $\mathcal{V}_{k,r}$, along with their inferred labels, can be used to expand the original labeled nodes set $\mathcal{V}_{k,L}$. The estimation of feature distribution could be augmented as:
$\boldsymbol{\mu}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}|}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}}\mathbf{x}^{prop}_{i},\quad\boldsymbol{s^{2}}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}|-1}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}}(\mathbf{x}^{prop}_{i}-\boldsymbol{\mu}_{k}^{c})^{2},$ (8)
where $\mathcal{V}_{k,r}^{c}$ represents the subset of reliable nodes in $\mathcal{V}_{k,r}$ belonging to class $c$. This augmentation ensures a more accurate and robust estimation of class-wise feature distributions by incorporating reliable nodes into the computation.
4.3 Global Pseudo-Graph Generation
After clients estimate the class-wise feature distribution statistics $\{\boldsymbol{\mu}_{k}^{c},\boldsymbol{s^{2}}_{k}^{c}\}$, they upload them along with sample quantity of each class $\{N_{k}^{c}\}$ to the server. The server aggregates them to recover the global class-wise feature distribution. Specifically, the global mean and global variance of class $c$ are computed unbiasedly as:
$N^{c}=\sum_{k=1}^{m}N_{k}^{c},\quad\boldsymbol{\mu}^{c}=\frac{1}{N^{c}}\sum_{k=1}^{m}N_{k}^{c}\boldsymbol{\mu}_{k}^{c},\quad\boldsymbol{s^{2}}^{c}=\frac{1}{N^{c}-m}(\sum_{k=1}^{m}(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c}+\sum_{k=1}^{m}N_{k}^{c}(\boldsymbol{\mu}_{k}^{c}-\boldsymbol{\mu}^{c})^{2}),$ (9)
which are used to guide the global pseudo-graph generation.
Inspired by the graph condensation [35, 51], we generate a small-size global pseudo-graph $G^{\prime}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ by aligning its feature distribution with the aggregated global distribution. This pseudo-graph serves as a representative of the overall graph data across all clients. The generation process is illustrated in the top of Figure 2. We first pre-set the number of nodes of each class in the pseudo-graph and initialize the trainable node features $\mathbf{X}^{\prime}$ from Gaussian noise. The adjacency matrix $\mathbf{A}^{\prime}$ is derived using a trainable link predictor $g_{\theta}$:
$\mathbf{A}^{\prime}_{i,j}=g_{\theta}(\mathbf{X}^{\prime},\delta)=\mathbb{I}(\delta\leq\sigma(\frac{g_{\theta}(\mathbf{x}_{i}\|\mathbf{x}_{j})+g_{\theta}(\mathbf{x}_{j}\|\mathbf{x}_{i})}{2})),$ (10)
[…]
Following [51], to optimize $\mathbf{X}^{\prime}$ and $g_{\theta}$, we first propagate node features $\mathbf{X}^{\prime}$ as in Eq. 5. Then we calculate the class-wise sample mean $\boldsymbol{\mu}^{\prime c}$ and sample variance $\boldsymbol{s^{2}}^{\prime c}$ of the propagated node features $\mathbf{X}^{\prime prop}$ following Eq. 6. Then we calculate the alignment loss:
$L_{align}=\sum_{c=0}^{C-1}\lambda_{c}((\boldsymbol{\mu}^{\prime c}-\boldsymbol{\mu}^{c})^{2}+(\boldsymbol{s^{2}}^{\prime c}-\boldsymbol{s^{2}}^{c})^{2}),$ (11)
where $\lambda_{c}$ is the proportion of nodes belonging to class $c$ relative to the total number of nodes. To ensure the smoothness of the pseudo-graph, smoothness loss is applied:
$L_{smt}=\frac{1}{\sum_{i,j}\mathbf{A}^{\prime}_{i,j}}\sum_{i,j}\mathbf{A}^{\prime}_{i,j}\exp{(-\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{2})}.$ (12)
We could optimize $\mathbf{X}^{\prime}$ and $g_{\theta}$ to generate the global pseudo-graph by the overall optimization objective: $\min_{\mathbf{X}^{\prime},\theta}\ (L_{align}+\alpha L_{smt})$. Once the optimization finishes, the server distributes $G^{\prime}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ to clients for further personalized training.
Privacy-preservation with Secure Aggregation. We demonstrate compatibility with Secure Aggregation protocols in uploading and aggregating $\{\boldsymbol{\mu}_{k}^{c},\boldsymbol{s^{2}}_{k}^{c},N_{k}^{c}\}$. Secure Aggregation is a crucial technique in federated learning that enables the server to compute the sum of large, user-held data vectors securely, without accessing individual client contributions. This property aligns well with the aggregation process in Eq. 9, as it can be expressed as a series of weighted average operations. The calculations of $N^{c}$ and $\boldsymbol{\mu}^{c}$ are inherently weighted average operations. The calculation of $\boldsymbol{s^{2}}^{c}$, while more complex, can also be decomposed into weighted average operations:
$\boldsymbol{s^{2}}^{c}=\frac{\sum_{k}^{m}(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c}}{N^{c}-m}+\frac{\sum_{k}^{m}N_{k}^{c}(\boldsymbol{\mu}_{k}^{c})^{2}-2\boldsymbol{\mu}^{c}\sum_{k}^{m}\boldsymbol{\mu}_{k}^{c}+N^{c}\sum_{k}^{m}(\boldsymbol{\mu}^{c})^{2}}{N^{c}-m}.$ (13)
Breaking this down further, the calculation involves the weighted averages of $(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c},N_{k}^{c}(\boldsymbol{\mu}_{k}^{c})^{2}$ and $\boldsymbol{\mu}_{k}^{c}$. Thus, the entire aggregation process for $\{\boldsymbol{\mu}_{k}^{c},\boldsymbol{s^{2}}_{k}^{c},N_{k}^{c}\}$ can be conducted without revealing individual client data contributions, making it fully compatible with Secure Aggregation.
4.4 Personalized Training with Node Adaptive Distillation
With the downloaded global pseudo-graph, each client trains its personalized model locally. However, the imbalanced local graph data often results in biased models that perform well on major classes but neglect minor classes. To achieve better personalization ($\uparrow$) and generalization (F1$\uparrow$), we propose a two-stage personalized training approach with node-adaptive distillation to leverage both global information in the global pseudo-graph and personal information in the local graph.
Stage 1
Each client trains a generalized model $M_{G}$ using only the global pseudo-graph, ensuring $M_{G}$ to capture global, balanced knowledge. The training objective is to minimize:
$L_{1^{st}}=L_{ce}(M_{G}(\mathbf{A}^{\prime},\mathbf{X}^{\prime}),\mathbf{Y}^{\prime}).$ (14)
A copy of $M_{G}$ is detached as a fixed teacher model for the next stage. $L_{ce}$ denotes the cross-entropy loss. Note that if the server knows the model architecture, an alternative is to conduct this stage on the server. However, we can adapt to more rigorous scenarios where models’ intellectual property needs to be protected. In those cases, it’s necessary to conduct this stage on each client.
Stage 2
Client $k$ fine-tunes $M_{G}$ with local graph data to derive a personalized model $M_{k}$. However, fine-tuning on imbalanced local graph data often leads to over-fitting on major classes while forgetting minor classes. As illustrated in Figure 5, the fine-tuned model shows improvement in major classes but suffers degradation in minor classes.
To mitigate this, we propose to utilize the global knowledge of $M_{G}$ and personalized knowledge in local graph data to train a better $M_{k}$ with the node adaptive distillation. For nodes of major classes, the model directly learns through supervised fine-tuning. For nodes of minor classes, $M_{k}$ additionally distills the predictive ability from $M_{G}$. The training objective for the second stage is to minimize:
$L_{2^{nd}}=L_{ft}+L_{dist},$ (15)
where $L_{ft}$ is the cross-entropy loss for supervised fine-tuning on local training data, and $L_{dist}$ is the weighted sum of point-wise KL divergence loss to distill global knowledge from $M_{G}$ into $M_{k}$:
$L_{dist}=\sum_{\mathbf{x}_{i}\in\mathbf{X}_{k}}\gamma_{i}L_{kl}(M_{k}(\mathbf{A}_{k},\mathbf{x}_{i}),M_{G}(\mathbf{A}_{k},\mathbf{x}_{i})),$ (16)
where $\gamma_{i}$ is a node-specific weighting factor to control the balance between local fine-tuning and global knowledge distillation. To determine $\gamma_{i}$ for each node, we re-use the soft labels $\tilde{\mathbf{y}}_{i}$ obtained in Section 4.2 and $\mathbf{w}_{dist}$ in Eq. 3:
$\gamma_{i}=\beta\tilde{\mathbf{y}}_{i}\cdot\mathbf{w}_{dist},$ (17)
where $\beta$ is a scaling hyper-parameter. Intuitively explained, if a node is predicted to class $c$ with a small $w_{dist}(c)$ (e.g. node $b$ in Figure 7), it likely belongs to major classes with high homophily. In this case, $\gamma_{i}$ is small, and $M_{i}$ primarily learns through supervised fine-tuning; If a node is predicted to class $c$ with a large $w_{dist}(c)$ or has a blended prediction (e.g. node $a$ in Figure 7), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus $\gamma_{i}$ should be large to incorporate the corresponding global knowledge from $M_{G}$.
Table 1: Performance on datasets under Louvain partition with 10 clients.
Method Cora CiteSeer PubMed ogbn-arxiv ogbn-products
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 67.17 $\pm$0.38 41.79 $\pm$0.52 59.82 $\pm$0.12 42.70 $\pm$0.34 81.74 $\pm$0.06 64.61 $\pm$0.18 66.46 $\pm$0.24 16.59 $\pm$0.57 85.25 $\pm$0.02 22.35 $\pm$0.19
FedAvg 69.68 $\pm$0.48 45.10 $\pm$0.75 63.92 $\pm$0.10 45.97 $\pm$0.43 78.24 $\pm$0.08 42.24 $\pm$0.13 67.13 $\pm$0.14 18.46 $\pm$0.23 85.34 $\pm$0.04 22.80 $\pm$0.04
FedPUB 68.50 $\pm$0.38 43.35 $\pm$0.27 61.80 $\pm$0.37 43.96 $\pm$1.05 81.63 $\pm$0.54 63.39 $\pm$0.69 65.34 $\pm$0.28 12.50 $\pm$0.20 83.91 $\pm$0.31 17.29 $\pm$0.52
FedGTA 42.61 $\pm$0.76 23.47 $\pm$1.06 67.12 $\pm$1.58 52.68 $\pm$1.81 75.99 $\pm$1.21 57.48 $\pm$1.44 58.30 $\pm$0.30 12.22 $\pm$0.26 75.87 $\pm$0.04 16.58 $\pm$0.15
FedTAD 70.10 $\pm$0.22 45.57 $\pm$0.27 63.93 $\pm$0.27 44.59 $\pm$0.41 78.21 $\pm$0.31 42.45 $\pm$0.80 68.77 $\pm$0.12 27.33 $\pm$0.50 OOM
FedSpray 63.14 $\pm$1.73 40.82 $\pm$1.54 54.89 $\pm$4.24 39.56 $\pm$3.58 77.89 $\pm$1.67 64.23 $\pm$2.03 68.68 $\pm$0.05 19.90 $\pm$0.16 84.06 $\pm$0.11 20.80 $\pm$0.05
FedLoG 66.24 $\pm$0.60 38.68 $\pm$0.54 61.19 $\pm$0.33 39.56 $\pm$3.58 79.40 $\pm$0.64 57.33 $\pm$0.47 OOM OOM
DENSE 69.89 $\pm$0.30 45.35 $\pm$0.55 64.09 $\pm$0.48 46.00 $\pm$0.39 78.16 $\pm$0.06 41.96 $\pm$0.33 68.80 $\pm$0.04 27.60 $\pm$0.36 86.12 $\pm$0.03 26.58 $\pm$0.06
Co-Boost 69.95 $\pm$0.40 45.43 $\pm$0.52 64.21 $\pm$0.27 46.00 $\pm$0.11 78.13 $\pm$0.05 41.84 $\pm$0.22 66.27 $\pm$0.08 16.37 $\pm$0.14 85.24 $\pm$0.00 22.31 $\pm$0.11
O-pFGL (Ours) 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34 70.93 $\pm$0.59 36.66 $\pm$0.22 86.63 $\pm$0.02 27.21 $\pm$0.15
5 Experiments
5.1 Experimental Setup
Datasets.
We conduct experiments on 12 datasets, including both transductive learning datasets: Cora, CiteSeer, PubMed [56], ogbn-arxiv, ogbn-products [23], Computers, Photo, CS, Physics datasets [40], and inductive learning datasets: Flickr [59], Reddit [19], and Reddit2 [59]. Table 4 summarizes these datasets. We employ two advanced graph partitions to simulate distributed subgraphs: Louvain-based Label Imbalance Split and Metis-based Label Imbalance Split [34].
Baselines.
We compare our method with 9 state-of-the-art methods, including 5 pFGL methods [^1] We do not include FedSage [63] and AdaFGL [32] since they need extra communication rounds to train neighborhood generators or federated optimization. They cannot train models within one-shot communication.: (1) FedPUB [2], (2) FedGTA [33], (3) FedTAD [69], (4) FedSpray [15], (5) FedLoG [30], and 2 OFL methods: (6) DENSE [62], (7) Co-Boost [9]. We also include (8) Standalone, which locally trains models, and (9) FedAvg with fine-tuning. Detailed descriptions are in the Appendix A.7.
Configurations.
We limit the communication round to a single round. We employ the two-layer GCN [31] with the hidden dimension set to 64. We employ both accuracy(%) and F1-macro(%) metrics. F1-macro averages the F1 scores across all classes, providing a more robust measurement of the model’s generalization, particularly under imbalanced data. We report the mean and standard deviation across three runs. Detailed configurations are described in the Appendix A.2.
Figure 3: Ablation study under Louvain partition.
Figure 4: Ablation study under Metis partition.
Figure 5: Client-level analysis on CiteSeer dataset. Left: Comparison of different personalized training methods. $M_{G}$ is the generalized model trained in Stage 1. Middle: The value of $H$ and $w_{dist}$ for each class. Right: The value of $H$ is determined by both samples quantity and homophily.
5.2 Experimental Results
5.2.1 Performance Comparison
The performance under the Louvain partition with 10 clients is shown in Table 1, and results with more clients and results under the Metis partition are shown in the Appendix A.3.1. Our method consistently outperforms on both accuracy and F1-macro, demonstrating the superiority of our method in deriving models with better personalization and local generalization. FedTAD encounters the OOM issue due to its graph diffusion operation, and FedLoG encounters the OOM issue due to its $k$-hop subgraph construction for each node. FedSpray and FedLoG also focus on minor and underrepresented classes, but they require numerous communications for optimization under non-IID scenarios, offering little advantage in one-shot communication. Furthermore, our method (1 communication round) even outperforms other multi-round federated learning methods (100 communication rounds) on average, which is shown in the Appendix A.3.2. More results on Computers, Photo, CS, and Physics datasets are listed in the Appendix A.3.6. We also include the experiments with other prevalent graph models (GraphSage [19], GAT [45], and SGC [49]) in Appendix A.3.7.
5.2.2 Ablation Study
We conduct the ablation study to analyze our method. We compare our method with 4 variants: (1) ft: only fine-tuning $M_{G}$ to obtain $M_{i}$ on local graph data, (2) ft(focal): fine-tuning with focal loss [39] to handle data imbalance, (3) ft(fixed $\gamma$): in the second stage, performing distillation with a fixed $\gamma_{i}$ for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE, and (5) ours: performing the complete method (including HRE).
The results under the Louvain partition and Metis partition are shown in Figure 4 and Figure 4, respectively. The focal loss cannot effectively address the non-IID graph data since it’s hard to determine the proper weights. Fixed $\gamma$ cannot balance improvements in accuracy and F1-macro as it cannot distinguish the majority from the minority. For nodes in major classes, the global knowledge of $M_{G}$ may have a negative impact, as demonstrated by the left sub-figure in Figure 5. This motivates us to develop node-adaptive distillation, which is more flexible to determine when to introduce global knowledge of $M_{G}$ in fine-tuning. HRE strategy further improves performance by generating a better global pseudo-graph with more precise estimations. Furthermore, we justify the rationale for considering homophily in determining the major classes by $H$ values in Eq. 2 in the right panel of Figure 5. The detailed analysis is in the Appendix A.3.3.
Table 2: Performance of methods on inductive datasets.
Method Flickr Reddit Reddit2
Acc. F1 Acc. F1 Acc. F1
Standalone 44.97 $\pm$0.70 12.84 $\pm$0.06 90.35 $\pm$0.02 28.79 $\pm$0.07 91.92 $\pm$0.01 32.53 $\pm$0.20
FedAvg 45.50 $\pm$0.97 13.66 $\pm$0.20 80.75 $\pm$0.81 16.85 $\pm$0.49 82.86 $\pm$0.62 19.20 $\pm$0.44
FedPUB 42.17 $\pm$0.00 8.32 $\pm$0.00 90.74 $\pm$0.02 27.86 $\pm$0.01 92.03 $\pm$0.02 30.37 $\pm$0.09
FedGTA 29.99 $\pm$1.71 14.38 $\pm$1.15 90.35 $\pm$0.02 28.76 $\pm$0.24 91.91 $\pm$0.38 32.71 $\pm$0.11
FedTAD 45.44 $\pm$0.47 13.85 $\pm$0.50 OOM OOM
FedSpray 43.08 $\pm$0.32 13.08 $\pm$1.75 89.07 $\pm$0.21 27.28 $\pm$0.24 91.28 $\pm$0.11 31.90 $\pm$0.17
FedLoG OOM OOM OOM
DENSE 44.60 $\pm$0.28 12.47 $\pm$0.39 90.32 $\pm$0.04 30.08 $\pm$0.10 92.00 $\pm$0.05 34.42 $\pm$0.29
Co-Boost 45.79 $\pm$0.54 13.52 $\pm$0.52 90.35 $\pm$0.03 28.62 $\pm$0.05 91.94 $\pm$0.02 32.68 $\pm$0.04
O-pFGL (Ours) 46.83 $\pm$0.44 14.31 $\pm$0.04 90.74 $\pm$0.11 30.09 $\pm$0.49 92.14 $\pm$0.05 35.51 $\pm$0.15
Table 3: Performance with heterogeneous models.
Method Cora CiteSeer ogbn-arxiv
Acc. F1 Acc. F1 Acc. F1
Standalone 68.37 $\pm$0.19 43.51 $\pm$0.31 62.50 $\pm$0.15 43.94 $\pm$0.09 66.13 $\pm$0.08 19.60 $\pm$0.38
FedTAD 68.18 $\pm$0.22 42.91 $\pm$0.34 61.81 $\pm$0.18 43.78 $\pm$0.14 69.08 $\pm$0.08 29.29 $\pm$0.44
FedSpray 65.73 $\pm$1.05 42.23 $\pm$0.34 53.14 $\pm$1.19 39.13 $\pm$1.17 68.63 $\pm$0.03 22.25 $\pm$0.26
DENSE 68.15 $\pm$0.47 43.02 $\pm$0.36 62.31 $\pm$0.51 44.27 $\pm$0.44 69.03 $\pm$0.02 29.64 $\pm$0.14
Co-Boost 68.38 $\pm$0.40 43.13 $\pm$0.52 62.22 $\pm$0.18 44.23 $\pm$0.05 69.08 $\pm$0.13 29.88 $\pm$0.27
O-pFGL (Ours) 75.04 $\pm$0.50 57.26 $\pm$1.93 70.87 $\pm$0.21 56.67 $\pm$0.20 70.98 $\pm$0.09 36.41 $\pm$0.43
Figure 6: Communication costs.
5.2.3 More Experiments
Inductive Performance.
The results on inductive datasets under the Louvain partition are shown in Table 2. More results are shown in the Appendix A.3.5. Our method is scalable and outperforms.
Model Heterogeneity.
We conduct experiments on 10 clients with heterogeneous models and configurations are described in Appendix A.2. Under model heterogeneity, most pFGL methods do not work since they rely on the model parameter aggregation. The results under Louvain and Metis partitions are presented in Table 3 and Table 11, respectively. The results demonstrate the superior performance of our method in both accuracy and F1-macro. Details are in the Appendix A.3.4.
Hyper-parameters Analysis.
We study our method’s sensitivity to the hyper-parameters in Appendix A.3.8. Our method is generally robust under a proper range of hyper-parameters values.
5.2.4 Communication Analysis
We analyze the communication costs both theoretically and empirically. Our method involves uploading the class-wise statistics ($O(Chf)$) and downloading the generated global pseudo-graph ($O(n^{\prime}f+n^{\prime 2})$, $n^{\prime}$ is the pre-set number of nodes in the global pseudo-graph, which is small as described in Appendix A.2). Both are independent of the model size. Figure 6 reports the communication costs of baseline methods and our method on the Cora datasets under the Louvain and Metis partition. The results show that our method has lower communication costs. We include the detailed theoretical analysis of communication, computation, and memory complexity in the Appendix A.4. We also introduce an equivalent server-efficient variant of our method in the Appendix A.5.
6 Conclusion
We propose the first one-shot personalized federated graph learning method that supports model heterogeneity and Secure Aggregation. Specifically, our method first estimates class-wise feature distribution statistics on each client, augmented by our proposed HRE, and then aggregates these statistics on the server to generate the global pseudo-graph. To further enhance personalization and generalization, we propose a novel two-stage personalized training with node-adaptive distillation, which balances local information and global insights. Extensive experiments demonstrate the superior performance of our method across various settings.
Appendix A Appendix
A.1 Illustration for Determining $\gamma_{i}$ in Node Adaptive Distillation
We illustrate how to determine $\gamma_{i}$ for each node in each client’s local graph. We pre-compute the vector $\mathbf{w}_{dist}\in\mathbb{R}^{C}$ by Eq. 3 and soft labels by Label Propagation in Section 4.2. Considering soft label of each node, if a node is predicted to class $c$ with a small $w_{dist}(c)$ (e.g. node $b$ in Figure 7), it likely belongs to major classes with high homophily. In this case, $\gamma_{i}$ is small, and $M_{i}$ primarily learns through supervised fine-tuning; If a node is predicted to class $c$ with a large $w_{dist}(c)$ or has a blended prediction (e.g. node $a$ in Figure 7), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus $\gamma_{i}$ should be large to incorporate the corresponding global knowledge from $M_{G}$.
Figure 7: The process to determine $\gamma_{i}$ for each node in node adaptive distillation in the second stage.
A.2 Experimental Setup
Table 4: The statistics and description of the datasets.
Dataset Task #Nodes #Edges #Features #Classes Train/Val/Test Description
Cora Transductive 2,708 10,556 1433 7 20%/40%/40% Citation Network
CiteSeer Transductive 3,327 4,732 3703 6 20%/40%/40% Citation Network
PubMed Transductive 19,717 44,338 500 3 20%/40%/40% Citation Network
Ogbn-arxiv Transductive 169,343 1,166,243 128 40 60%/20%/20% Citation Network
Ogbn-products Transductive 2,449,029 61,859,140 100 47 10%/5%/85% Co-purchase Network
Computers Transductive 13,381 245,778 767 10 20%/40%/40% Co-purchase Network
Photo Transductive 7,487 119,043 745 8 20%/40%/40% Co-purchase Network
CS Transductive 18,333 81,894 6,805 15 20%/40%/40% Co-author Network
Physics Transductive 34,493 247,962 8,415 5 20%/40%/40% Co-author Network
Flickr Inductive 89,250 899,756 500 7 60%/20%/20% Image Network
Reddit Inductive 232,965 114,615,892 602 41 80%/10%/10% Social Network
Reddit2 Inductive 232,965 23,213,838 602 41 65%/10%/25% Social Network
Dataset Description.
We conduct experiments on 12 graph datasets, and the summary is listed in Table 4. Our selected datasets cover a wide range from 1) small-scale to large-scale, 2) transductive to inductive, and 3) academic domain to industry domain.
Hyper-parameters and Model Configurations.
In our one-shot personalized federated graph learning setting, we limit the communication round of the 5 pFGL methods and the FedAvg method to a single round. For 2 OFL methods (DENSE and Co-Boost), the generators are trained to generate node features, and the topology structure is constructed using the $K$-Nearest Neighbors strategy as outlined in [69]. Note that we perform additional fine-tuning processes on FedAvg, FedTAD, DENSE, and Co-Boost to ensure better performance.
For each client, we employ a two-layer GCN [31] with the hidden layer dimension set to 64 by default. For baseline methods, the number of local training epochs is tuned from 1 to 100, and hyper-parameters are set as recommended in their respective papers or fine-tuned for optimal overall performance. For our method, we keep the size of the generated global pseudo-graph as small as possible. On small-scale Cora, CiteSeer, and PubMed datasets, we set the number of nodes in each class in the global pseudo-graph to 1. For other datasets, we set the number of nodes in each class to $p$ percent of the total number of nodes of the corresponding class in the training set. We set $p$ as 0.25%, 0.04%, 1%, 1%, 1%, 0.5%, 0.2%, 0.05%, 0.05% on ogbn-arxiv, ogbn-products, Computers, Photo, CS, Physics, Flickr, Reddit, and Reddit2 datasets. We use the Adam optimizer and tune the learning rate. We tune $f_{th}$ within [0.95, 1] and tune $d_{th}$ according to the scale of the graphs. We tune the $\beta$ within [0.1, 1] to control the range of $\gamma_{i}$. For other hyper-parameters in global pseudo-graph generation, we adopt the setting in [51].
Under model heterogeneity, we adopt 10 different GNN models categorized into two types: coupled GNNs and decoupled GNNs. For coupled GNNs, we adopt 6 models, which are: 2-layer GCN with 64, 128, 256, and 512 hidden dimensions, and 3-layer GCN with 64 and 512 hidden dimensions. For decoupled GNNs, we adopt APPNP [16], SSGC [68], SGC ($K$=2) and SGC ($K$=4) [49].
For metrics, we employ both accuracy(%) and F1-macro(%) metrics to comprehensively evaluate the performance. F1-macro averages the F1 scores across all classes, providing a more robust measurement of the model’s generalization, particularly under imbalanced data. For each experiment, we report the mean and standard deviation across three runs.
Experimental Environments.
The experimental machine is with Intel(R) Xeon(R) CPU E5-2680 v4@2.40GHz, and 6$\times$×NVIDIA GeForce RTX 3090. The operating system is Ubuntu 20.04, and the version of PyTorch is 2.1.2. We use PyG 2.5.2 [11].
Table 5: Performance of methods on datasets under Louvain partition with 20 clients.
Method Cora CiteSeer PubMed ogbn-arxiv ogbn-products
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 61.57 $\pm$0.84 30.35 $\pm$0.59 57.86 $\pm$0.62 36.44 $\pm$0.52 81.84 $\pm$0.04 57.95 $\pm$0.62 66.56 $\pm$0.16 13.53 $\pm$0.34 86.44 $\pm$0.01 17.34 $\pm$0.05
FedAvg 63.21 $\pm$0.45 31.77 $\pm$0.64 61.51 $\pm$0.27 40.60 $\pm$0.40 78.59 $\pm$0.14 41.60 $\pm$0.28 67.17 $\pm$0.12 15.33 $\pm$0.44 86.48 $\pm$0.01 17.57 $\pm$0.12
FedPUB 62.52 $\pm$0.27 30.96 $\pm$0.25 59.40 $\pm$0.29 37.59 $\pm$0.19 81.41 $\pm$0.14 56.81 $\pm$1.94 65.83 $\pm$0.33 10.32 $\pm$0.53 83.47 $\pm$1.18 11.23 $\pm$0.74
FedGTA 37.03 $\pm$1.34 14.74 $\pm$0.76 62.02 $\pm$0.83 42.78 $\pm$0.92 64.01 $\pm$0.98 36.47 $\pm$1.33 57.59 $\pm$1.39 7.72 $\pm$0.67 79.04 $\pm$0.00 12.15 $\pm$0.10
FedTAD 63.16 $\pm$0.44 31.72 $\pm$0.52 62.32 $\pm$0.46 41.04 $\pm$0.55 78.64 $\pm$0.11 41.66 $\pm$0.16 68.90 $\pm$0.14 23.31 $\pm$0.22 OOM
FedSpray 60.15 $\pm$1.05 29.30 $\pm$0.61 55.43 $\pm$0.79 36.77 $\pm$0.63 80.28 $\pm$0.30 49.29 $\pm$0.73 68.88 $\pm$0.14 17.56 $\pm$0.27 85.43 $\pm$0.09 16.08 $\pm$0.10
FedLoG 60.92 $\pm$0.28 28.68 $\pm$0.12 58.29 $\pm$0.15 33.99 $\pm$0.34 79.73 $\pm$0.23 56.26 $\pm$0.52 OOM OOM
DENSE 63.24 $\pm$0.32 31.74 $\pm$0.44 61.71 $\pm$0.21 40.69 $\pm$0.44 78.59 $\pm$0.15 41.63 $\pm$0.31 68.81 $\pm$0.04 22.97 $\pm$0.18 86.95 $\pm$0.01 21.14 $\pm$0.00
Co-Boost 63.21 $\pm$0.41 31.70 $\pm$0.49 61.89 $\pm$0.42 40.83 $\pm$0.54 78.67 $\pm$0.10 41.76 $\pm$0.27 66.77 $\pm$0.11 13.90 $\pm$0.14 86.43 $\pm$0.01 17.35 $\pm$0.14
O-pFGL (Ours) 69.84 $\pm$1.76 49.37 $\pm$1.98 68.69 $\pm$0.04 50.26 $\pm$0.80 82.04 $\pm$0.20 58.79 $\pm$0.68 70.61 $\pm$0.05 32.28 $\pm$0.68 87.03 $\pm$0.00 21.87 $\pm$0.05
A.3 More Experimental Results and Detailed Analysis
A.3.1 Performance Comparison
We first conduct experiments in the transductive learning setting and report the accuracy and F1-macro metrics. The datasets range from small-scale graphs to large-scale graphs with millions of nodes. We also vary the number of clients to evaluate the stability of methods. With 10 clients, the experimental results under Louvain and Metis partitions are shown in Table 1 and Table 6, respectively. With 20 clients, the experimental results under Louvain and Metis partitions are shown in Table 5 and Table 8, respectively. Our method consistently outperforms others in terms of accuracy and F1-macro. Under the non-IID scenarios, pFGL methods show marginal or even negative performance gain over Standalone, highlighting the ineffectiveness of current pFGL methods in personalizing models on non-IID graph data in a single communication round. The state-of-the-art OFL methods (DENSE and Co-Boost) also have difficulty generating high-quality pseudo-data by the ensemble model when each model is trained on non-IID local graph data. FedTAD encounters OOM when handling million-scale graphs due to its graph diffusion operation. FedLoG encounters OOM issues due to its node-wise $k$-hop subgraph construction. In contrast, our method remains scalable. FedGTA performs badly because its aggregation is based on the unreliable local models’ predictions under non-IID data in a single round. FedSpray focuses on minor classes, but it requires numerous communications for optimization, offering little advantage over others in one-shot communication. FedLoG focuses on local generalization, but it cannot work well under severe non-IID scenarios within one-shot communication.
Table 6: Performance of methods on datasets under Metis partition with 10 clients.
Method Cora CiteSeer PubMed ogbn-arxiv ogbn-products
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 75.15 $\pm$0.08 31.00 $\pm$0.09 64.72 $\pm$0.37 37.36 $\pm$0.38 83.24 $\pm$0.09 56.61 $\pm$0.77 66.39 $\pm$0.26 24.05 $\pm$0.70 85.57 $\pm$0.02 23.54 $\pm$0.10
FedAvg 76.18 $\pm$0.14 31.42 $\pm$0.08 66.51 $\pm$0.34 35.89 $\pm$0.54 79.26 $\pm$0.09 33.75 $\pm$0.18 67.04 $\pm$0.13 26.06 $\pm$0.66 85.63 $\pm$0.01 24.16 $\pm$0.27
FedPUB 75.31 $\pm$0.20 32.00 $\pm$0.53 64.94 $\pm$0.42 37.50 $\pm$0.57 82.69 $\pm$0.10 56.44 $\pm$1.21 64.86 $\pm$0.62 18.29 $\pm$1.14 84.43 $\pm$0.56 17.24 $\pm$0.64
FedGTA 53.85 $\pm$0.29 11.74 $\pm$0.66 64.03 $\pm$0.28 31.33 $\pm$0.54 78.57 $\pm$0.58 46.47 $\pm$0.94 55.90 $\pm$1.23 12.72 $\pm$1.53 74.36 $\pm$0.34 16.67 $\pm$0.19
FedTAD 76.36 $\pm$0.26 31.99 $\pm$0.22 66.76 $\pm$0.26 36.06 $\pm$0.42 79.41 $\pm$0.07 34.19 $\pm$0.13 68.96 $\pm$0.28 34.91 $\pm$0.20 OOM
FedSpray 74.31 $\pm$0.34 31.95 $\pm$0.69 62.49 $\pm$0.85 36.04 $\pm$0.54 81.02 $\pm$0.56 60.97 $\pm$1.27 68.21 $\pm$0.18 26.57 $\pm$0.17 86.45 $\pm$0.05 22.02 $\pm$0.06
FedLoG 75.08 $\pm$0.21 23.75 $\pm$0.35 64.92 $\pm$0.36 33.72 $\pm$0.58 75.52 $\pm$6.73 48.26 $\pm$4.12 OOM OOM
DENSE 76.21 $\pm$0.08 31.64 $\pm$0.24 65.62 $\pm$0.27 37.88 $\pm$0.24 79.29 $\pm$0.05 33.81 $\pm$0.09 68.89 $\pm$0.04 34.34 $\pm$0.34 86.43 $\pm$0.03 28.42 $\pm$0.25
Co-Boost 76.54 $\pm$0.31 31.90 $\pm$0.36 64.55 $\pm$0.73 31.37 $\pm$0.75 76.56 $\pm$0.36 29.89 $\pm$0.34 66.22 $\pm$0.02 23.58 $\pm$0.18 85.59 $\pm$0.03 23.63 $\pm$0.25
O-pFGL (Ours) 81.79 $\pm$0.23 50.85 $\pm$1.27 72.76 $\pm$0.24 50.94 $\pm$0.80 83.99 $\pm$0.18 61.03 $\pm$1.21 71.06 $\pm$0.09 41.35 $\pm$0.37 86.90 $\pm$0.01 28.90 $\pm$0.05
In addition to improving accuracy, our method consistently achieves a significant increase in F1-macro, an often overlooked metric in previous studies. This improvement highlights that our method enhances both personalization and generalization, resulting in more robust models.
We also compare our method with ReNode [6], a representative graph imbalance learning method, which aims to train robust graph models under imbalanced graph data. The comparison results are shown in Table 7. The results show that ReNode cannot effectively model the imbalanced graph data in our scenarios.
Table 7: Comparison with the graph imbalance learning method.
Louvain Cora CiteSeer PubMed
Acc. F1 Acc. F1 Acc. F1
ReNode 68.07 $\pm$0.29 41.70 $\pm$0.57 61.57 $\pm$0.25 41.91 $\pm$0.26 74.64 $\pm$0.36 35.59 $\pm$0.69
O-pFGL (Ours) 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34
Metis Cora CiteSeer PubMed
Acc. F1 Acc. F1 Acc. F1
ReNode 76.44 $\pm$0.06 27.69 $\pm$0.45 65.89 $\pm$0.23 37.74 $\pm$0.15 78.81 $\pm$0.14 32.94 $\pm$0.18
O-pFGL (Ours) 81.79 $\pm$0.23 50.85 $\pm$1.27 72.76 $\pm$0.24 50.94 $\pm$0.80 83.99 $\pm$0.18 61.03 $\pm$1.21
Table 8: Performance of methods on datasets under Metis partition with 20 clients.
Method Cora CiteSeer PubMed ogbn-arxiv ogbn-products
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 71.57 $\pm$0.29 20.89 $\pm$0.06 61.36 $\pm$0.27 26.19 $\pm$0.40 83.24 $\pm$0.05 55.21 $\pm$0.35 66.77 $\pm$0.17 19.76 $\pm$0.49 86.44 $\pm$0.00 18.32 $\pm$0.14
FedAvg 73.14 $\pm$0.22 18.38 $\pm$0.39 62.62 $\pm$0.48 28.48 $\pm$0.38 79.86 $\pm$0.09 34.81 $\pm$0.05 67.54 $\pm$0.10 21.68 $\pm$0.16 86.52 $\pm$0.00 19.04 $\pm$0.09
FedPUB 72.32 $\pm$0.32 20.16 $\pm$0.82 60.73 $\pm$0.14 29.40 $\pm$0.16 82.98 $\pm$0.21 50.51 $\pm$0.26 65.41 $\pm$0.34 14.45 $\pm$0.59 83.94 $\pm$1.18 11.66 $\pm$0.60
FedGTA 55.04 $\pm$0.14 10.60 $\pm$0.07 59.43 $\pm$0.40 20.04 $\pm$0.82 74.34 $\pm$0.36 35.18 $\pm$0.61 58.76 $\pm$0.74 10.77 $\pm$0.44 78.35 $\pm$0.06 13.81 $\pm$0.26
FedTAD 73.55 $\pm$0.37 18.78 $\pm$0.31 62.60 $\pm$0.40 28.44 $\pm$0.20 77.42 $\pm$0.47 31.17 $\pm$0.52 69.29 $\pm$0.07 29.84 $\pm$0.34 OOM
FedSpray 71.44 $\pm$0.04 21.49 $\pm$0.15 58.98 $\pm$1.87 29.35 $\pm$1.52 81.05 $\pm$0.33 40.51 $\pm$0.56 68.86 $\pm$0.44 22.49 $\pm$0.44 85.28 $\pm$0.05 17.31 $\pm$0.04
FedLoG 73.79 $\pm$0.12 18.72 $\pm$0.30 61.66 $\pm$0.15 26.16 $\pm$0.35 80.35 $\pm$0.83 49.59 $\pm$0.64 OOM OOM
DENSE 72.62 $\pm$0.17 20.19 $\pm$0.13 61.36 $\pm$0.19 29.66 $\pm$0.20 79.94 $\pm$0.04 34.87 $\pm$0.07 69.21 $\pm$0.15 29.31 $\pm$0.54 86.97 $\pm$0.00 22.94 $\pm$0.12
Co-Boost 73.42 $\pm$0.06 18.45 $\pm$0.24 61.30 $\pm$0.56 25.87 $\pm$0.78 79.98 $\pm$0.06 35.04 $\pm$0.27 66.72 $\pm$0.06 19.48 $\pm$0.28 86.43 $\pm$0.01 18.51 $\pm$0.24
O-pFGL (Ours) 78.31 $\pm$0.88 37.21 $\pm$1.96 68.88 $\pm$0.22 41.95 $\pm$0.25 83.94 $\pm$0.05 56.61 $\pm$1.36 70.87 $\pm$0.18 36.06 $\pm$0.43 87.15 $\pm$0.02 23.41 $\pm$0.26
A.3.2 Compared with Multi-Rounds FL Methods
To demonstrate the superiority of our method, we further compare our method with multi-round federated learning methods, including FedAvg, FedPUB, FedGTA, FedTAD, FedSpray, and FedLoG. We set the communication round of these 6 methods to 100. The experimental results on 5 multi-scale graph datasets under Louvain and Metis partitions are shown in Table 9 and Table 10, respectively. Our method still outperforms in most cases and remains the best average performance.
Table 9: Comparison with FL methods with 100 communication rounds under Louvain partition.
Method Cora CiteSeer PubMed ogbn-arxiv ogbn-products Average
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
FedAvg 75.27 $\pm$0.17 61.74 $\pm$0.21 67.11 $\pm$0.48 51.86 $\pm$0.81 85.22 $\pm$0.14 71.17 $\pm$0.22 68.10 $\pm$0.07 30.72 $\pm$0.14 86.58 $\pm$0.02 28.28 $\pm$0.07 76.46 48.75
FedPUB 78.32 $\pm$0.37 64.59 $\pm$0.80 69.00 $\pm$0.74 56.43 $\pm$0.68 83.12 $\pm$0.08 70.18 $\pm$0.36 65.34 $\pm$0.28 12.50 $\pm$0.20 85.18 $\pm$0.14 16.31 $\pm$0.46 76.19 44.00
FedGTA 73.32 $\pm$0.17 60.08 $\pm$0.32 67.12 $\pm$1.58 52.68 $\pm$1.81 84.39 $\pm$0.08 72.06 $\pm$0.12 58.30 $\pm$0.30 12.22 $\pm$0.26 83.25 $\pm$0.17 24.71 $\pm$0.10 73.28 44.35
FedTAD 74.62 $\pm$0.44 60.37 $\pm$1.01 67.68 $\pm$0.39 53.43 $\pm$0.82 84.70 $\pm$0.21 68.56 $\pm$0.72 68.77 $\pm$0.12 27.33 $\pm$0.50 OOM N/A N/A
FedSpray 70.08 $\pm$0.41 56.46 $\pm$1.30 68.31 $\pm$0.88 55.07 $\pm$0.94 84.04 $\pm$0.10 74.06 $\pm$0.29 70.20 $\pm$0.08 34.09 $\pm$0.33 84.06 $\pm$0.11 20.80 $\pm$0.05 75.34 48.10
FedLoG 69.08 $\pm$1.96 53.80 $\pm$1.43 61.19 $\pm$0.33 39.42 $\pm$0.36 79.40 $\pm$0.62 57.33 $\pm$0.47 OOM OOM N/A N/A
O-pFGL (Ours) 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34 70.93 $\pm$0.59 36.66 $\pm$0.22 86.63 $\pm$0.02 27.21 $\pm$0.15 77.66 50.02
Table 10: Comparison with FL methods with 100 communication rounds under Metis partition.
Method Cora CiteSeer PubMed ogbn-arxiv ogbn-products Average
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
FedAvg 77.10 $\pm$0.30 45.59 $\pm$1.81 67.30 $\pm$0.30 42.88 $\pm$0.10 83.70 $\pm$0.03 66.65 $\pm$0.18 68.76 $\pm$0.17 36.77 $\pm$0.07 86.80 $\pm$0.01 30.02 $\pm$0.10 76.73 44.38
FedPUB 76.31 $\pm$0.61 37.22 $\pm$1.68 66.07 $\pm$1.05 40.65 $\pm$0.97 84.66 $\pm$1.04 68.19 $\pm$1.51 64.86 $\pm$0.62 18.29 $\pm$1.14 85.83 $\pm$0.35 19.45 $\pm$0.35 75.55 36.76
FedGTA 75.67 $\pm$0.52 44.91 $\pm$1.54 64.03 $\pm$0.28 31.33 $\pm$0.54 85.04 $\pm$0.06 65.87 $\pm$0.32 55.90 $\pm$1.23 12.72 $\pm$1.53 74.36 $\pm$0.34 16.67 $\pm$0.19 71.00 34.30
FedTAD 77.33 $\pm$0.44 45.49 $\pm$0.51 66.76 $\pm$0.26 36.06 $\pm$0.42 79.41 $\pm$0.07 34.19 $\pm$0.13 68.96 $\pm$0.28 34.91 $\pm$0.20 OOM N/A N/A
FedSpray 74.31 $\pm$0.34 31.95 $\pm$0.69 66.40 $\pm$0.68 49.06 $\pm$1.61 84.17 $\pm$0.14 70.45 $\pm$0.34 70.12 $\pm$0.15 41.26 $\pm$0.26 86.45 $\pm$0.05 22.02 $\pm$0.06 76.29 42.95
FedLoG 75.08 $\pm$0.21 23.75 $\pm$0.35 64.92 $\pm$0.36 33.72 $\pm$0.58 75.52 $\pm$6.74 48.26 $\pm$4.12 OOM OOM N/A N/A
O-pFGL (Ours) 81.79 $\pm$0.23 50.85 $\pm$1.27 72.76 $\pm$0.24 50.94 $\pm$0.80 83.99 $\pm$0.18 61.03 $\pm$1.21 71.06 $\pm$0.09 41.35 $\pm$0.37 86.90 $\pm$0.01 28.90 $\pm$0.05 79.30 46.61
A.3.3 Detailed Ablation Study
We conduct the ablation study to evaluate the performance gain of our method. We compare our method with 4 variants: (1) ft: only fine-tuning $M_{G}$ to obtain $M_{i}$ on local graph data, (2) ft(focal): fine-tuning with focal loss [39] to handle data imbalance, (3) ft(fixed $\gamma$): in the second stage, performing distillation with a fixed $\gamma$ for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE, and (5) ours: performing the complete method (including HRE).
Figure 4 and Figure 4 show our experimental results of accuracy and F1-macro under Louvain and Metis partition, respectively. We make the following 4 observations. (O1) Comparing ft and ft(focal), the focal loss in ft(focal) cannot effectively address the non-IID graph data. We believe the reasons are that it’s hard to determine the proper weights under extreme non-IID scenarios. (O2) Comparing ft and ft(fixed $\gamma$), a fixed $\gamma$ in distillation cannot balance the improvements of accuracy and F1-macro. We believe the reason is that it does not distinguish the majority and minority in the distillation. For nodes in major classes with high homophily, the global knowledge of $M_{G}$ may have a negative impact, as demonstrated by the left panel in Figure 5. (O3) Comparing ft and ours w.o. HRE, the node adaptive distillation is more flexible in determining when to introduce the global knowledge of $M_{G}$ in fine-tuning. Thus ours w.o. HRE achieves better accuracy and F1-macro. (O4) Comparing ours w.o. HRE and ours, the HRE strategy further improves performance by generating a higher-quality global pseudo-graph through more precise estimations.
We make a further ablation at the client level in Figure 5, evaluating different methods on a client using the CiteSeer dataset with the Louvain partition. As shown in the middle figure, class 3 and class 5 are regarded as major classes due to their larger quantity and relatively high homophily, while class 2 and class 1 are regarded as minor classes with fewer nodes and lower homophily. We make the following 3 observations. (O1) Compared to $M_{G}$, the ft method shows performance gain on major class 3 (+1.57%) and class 5 (+0.56%) but degrades performance on minor class 4 (-1.09%), class 2 (-2.78%), and class 1 (-3.22%). The performance differences after fine-tuning, shown in the left figure, align with the $H$ values in the middle figure, highlighting the limitations of fine-tuning. (O2) The ft(fixed $\gamma$) method cannot handle this limitation effectively. With a fixed $\gamma$, the performance on major class 5 worsens (-0.42%), and the performance on minor classes still degrades. This phenomenon motivates us to develop node-adaptive distillation. (O3) Our method, with Class-aware Distillation Factor $w_{dist}$ and node adaptive distillation, boosts performance on all classes by more than +2.4%.
We further justify the rationale for considering homophily in determining the major classes by $H$ values in Eq. 2. The right panel of Figure 5 shows that while class 1 and class 4 have slightly more nodes, their average homophily is low ($<$ 0.6), leading to less compact communities in the local graph. As a result, they are often neglected during fine-tuning, causing performance drops of -3.22% and -1.09%, respectively. In contrast, class 3 exhibits high average homophily ($>$ 0.7), forming compact connections, allowing it to dominate the fine-tuning process with a performance improvement of +1.57%.
A.3.4 Performance under Model Heterogeneity
We conduct experiments on 10 clients with heterogeneous models. The model configuration is detailed in Section A.2. Under model heterogeneity, most pFGL methods (except FedTAD and FedSpray) do not work since they rely on the model parameter aggregation. Thus, we compare our method with FedTAD, FedSpray, DENSE, and Co-Boost. The results under Louvain and Metis partitions are presented in Table 3 and Table 11. The results demonstrate the superior performance of our method in both accuracy and F1-macro.
Table 11: Performance of methods on 10 clients with heterogeneous models under Metis partition.
Method Cora CiteSeer ogbn-arxiv
Acc. F1 Acc. F1 Acc. F1
Standalone 74.33 $\pm$0.14 30.73 $\pm$0.18 64.53 $\pm$0.34 37.52 $\pm$0.33 65.18 $\pm$0.11 25.25 $\pm$0.14
FedTAD 74.37 $\pm$0.30 31.27 $\pm$0.25 64.75 $\pm$0.64 37.29 $\pm$0.34 68.57 $\pm$0.08 34.32 $\pm$0.44
FedSpray 73.09 $\pm$0.69 30.84 $\pm$0.65 60.38 $\pm$0.67 34.96 $\pm$0.16 67.39 $\pm$0.11 27.88 $\pm$0.24
DENSE 73.51 $\pm$0.99 30.62 $\pm$0.42 65.33 $\pm$0.24 37.62 $\pm$0.09 68.56 $\pm$0.03 34.64 $\pm$0.03
Co-Boost 73.19 $\pm$0.72 30.56 $\pm$0.33 65.40 $\pm$0.09 37.66 $\pm$0.19 68.55 $\pm$0.09 34.66 $\pm$0.21
O-pFGL (Ours) 79.64 $\pm$0.94 44.86 $\pm$2.04 71.82 $\pm$0.34 47.57 $\pm$0.98 70.31 $\pm$0.06 40.31 $\pm$0.11
Table 12: Performance on inductive datasets under Louvain partition with 20 clients.
Method Flickr Reddit Reddit2
Acc. F1 Acc. F1 Acc. F1
Standalone 42.10 $\pm$0.14 15.39 $\pm$0.25 89.90 $\pm$0.06 13.08 $\pm$0.21 91.09 $\pm$0.02 16.52 $\pm$0.04
FedAvg 28.86 $\pm$1.96 13.55 $\pm$0.36 83.99 $\pm$0.28 7.81 $\pm$0.08 85.94 $\pm$0.30 8.48 $\pm$0.06
FedPUB 42.06 $\pm$0.00 8.31 $\pm$0.00 90.34 $\pm$0.08 11.95 $\pm$0.27 91.23 $\pm$0.02 13.96 $\pm$0.11
FedGTA 40.78 $\pm$4.26 14.11 $\pm$1.51 89.83 $\pm$0.05 12.90 $\pm$0.32 91.14 $\pm$0.05 16.64 $\pm$0.12
FedTAD 41.61 $\pm$0.42 15.88 $\pm$0.31 OOM OOM
FedSpray 42.32 $\pm$1.20 14.03 $\pm$0.08 89.41 $\pm$0.06 10.71 $\pm$0.14 90.51 $\pm$0.05 15.29 $\pm$0.06
FedLoG OOM OOM OOM
DENSE 42.45 $\pm$0.72 11.67 $\pm$1.07 89.89 $\pm$0.05 14.27 $\pm$0.14 91.25 $\pm$0.03 18.14 $\pm$0.44
Co-Boost 42.07 $\pm$0.07 15.28 $\pm$0.45 89.93 $\pm$0.03 13.08 $\pm$0.19 91.12 $\pm$0.03 16.62 $\pm$0.15
O-pFGL (Ours) 44.00 $\pm$0.29 15.60 $\pm$0.47 90.38 $\pm$0.05 15.70 $\pm$0.47 91.34 $\pm$0.03 18.34 $\pm$0.18
Table 13: Performance on inductive datasets under Metis partition with 10 clients.
Method Flickr Reddit Reddit2
Acc. F1 Acc. F1 Acc. F1
Standalone 47.47 $\pm$0.12 12.89 $\pm$0.47 88.80 $\pm$0.09 42.34 $\pm$0.40 90.86 $\pm$0.04 44.78 $\pm$0.35
FedAvg 34.80 $\pm$3.92 13.45 $\pm$0.59 78.83 $\pm$0.07 24.21 $\pm$0.30 79.30 $\pm$0.18 26.26 $\pm$0.45
FedPUB 42.65 $\pm$0.40 10.09 $\pm$2.14 88.97 $\pm$0.18 41.01 $\pm$0.18 90.72 $\pm$0.07 41.66 $\pm$0.30
FedGTA 35.88 $\pm$6.10 13.04 $\pm$3.66 88.82 $\pm$0.11 42.02 $\pm$0.21 90.80 $\pm$0.07 44.59 $\pm$0.27
FedTAD 47.55 $\pm$0.08 13.07 $\pm$0.23 OOM OOM
FedSpray 43.77 $\pm$1.58 13.37 $\pm$0.42 87.71 $\pm$0.18 41.65 $\pm$0.29 89.99 $\pm$0.23 43.74 $\pm$0.26
FedLoG OOM OOM OOM
DENSE 45.26 $\pm$0.84 11.56 $\pm$0.84 88.72 $\pm$0.08 44.31 $\pm$0.23 90.01 $\pm$0.10 46.68 $\pm$0.03
Co-Boost 47.42 $\pm$0.10 12.90 $\pm$0.61 88.71 $\pm$0.07 42.08 $\pm$0.27 90.81 $\pm$0.06 44.56 $\pm$0.20
O-pFGL (Ours) 47.89 $\pm$0.11 13.11 $\pm$0.34 89.59 $\pm$0.10 45.31 $\pm$0.36 91.16 $\pm$0.04 48.56 $\pm$0.31
A.3.5 Inductive Performance
We conduct experiments on three inductive graph datasets: Flickr, Reddit, and Reddit2, and vary the number of clients. Since Flickr has lower homophily, we use the 2-layer SIGN [13] for better graph learning. With 10 clients, experimental results under Louvain and Metis partitions are shown in Table 2 and Table 13, respectively. With 20 clients, experimental results under Louvain and Metis partitions are shown in Table 12 and Table 14, respectively. Our method is scalable and generally performs best.
Table 14: Performance on inductive datasets under Metis partition with 20 clients.
Method Flickr Reddit Reddit2
Acc. F1 Acc. F1 Acc. F1
Standalone 47.10 $\pm$0.20 13.22 $\pm$0.37 87.52 $\pm$0.03 28.22 $\pm$0.10 89.59 $\pm$0.01 25.35 $\pm$0.08
FedAvg 33.44 $\pm$2.08 12.89 $\pm$0.45 80.71 $\pm$0.28 14.64 $\pm$0.48 82.90 $\pm$0.39 13.91 $\pm$0.41
FedPUB 42.10 $\pm$0.00 8.30 $\pm$0.00 87.36 $\pm$0.56 26.45 $\pm$0.71 89.89 $\pm$0.06 22.63 $\pm$0.31
FedGTA 38.75 $\pm$9.16 11.38 $\pm$3.46 87.49 $\pm$0.01 27.95 $\pm$0.22 89.55 $\pm$0.03 25.47 $\pm$0.29
FedTAD 47.17 $\pm$0.13 13.03 $\pm$0.38 OOM OOM
FedSpray 43.35 $\pm$0.96 14.00 $\pm$0.55 86.66 $\pm$0.09 27.60 $\pm$0.20 88.99 $\pm$0.01 24.23 $\pm$0.39
FedLoG OOM OOM OOM
DENSE 44.13 $\pm$0.97 11.45 $\pm$0.39 87.60 $\pm$0.02 30.49 $\pm$0.17 89.97 $\pm$0.05 27.72 $\pm$0.16
Co-Boost 47.08 $\pm$0.22 13.11 $\pm$0.66 87.52 $\pm$0.06 28.09 $\pm$0.12 89.67 $\pm$0.07 25.62 $\pm$0.22
O-pFGL (Ours) 47.08 $\pm$0.27 13.37 $\pm$0.24 88.28 $\pm$0.01 31.01 $\pm$0.47 90.25 $\pm$0.07 30.74 $\pm$0.25
A.3.6 More Datasets
We further provide more experimental results on Computers, Photo, CS, and Physics datasets. Computers and Photo are Amazon co-purchase networks and CS and Physics are Co-author networks. The results under Louvain and Metis partitions are shown in Figure 15 and Figure 16, respectively. Our method consistently outperforms others in terms of both accuracy and F1-macro, which demonstrates the general advantages of our method.
Table 15: Performance of methods on more datasets under the Louvain partition.
Method Computers Photo CS Physics
Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 87.34 $\pm$0.06 38.27 $\pm$0.91 89.16 $\pm$0.20 40.46 $\pm$1.31 87.18 $\pm$0.02 42.97 $\pm$0.07 93.65 $\pm$0.06 68.40 $\pm$0.53
FedAvg 87.18 $\pm$0.75 38.94 $\pm$0.51 89.26 $\pm$0.13 41.11 $\pm$0.77 87.10 $\pm$0.11 44.86 $\pm$0.23 93.55 $\pm$0.03 70.72 $\pm$0.38
FedPUB 83.51 $\pm$2.46 31.48 $\pm$2.64 89.43 $\pm$0.05 38.17 $\pm$1.24 87.01 $\pm$0.12 41.99 $\pm$0.38 93.55 $\pm$1.22 64.95 $\pm$1.91
FedGTA 56.73 $\pm$14.13 13.41 $\pm$3.08 53.68 $\pm$3.75 15.09 $\pm$1.37 83.26 $\pm$0.11 40.34 $\pm$1.10 91.81 $\pm$0.22 64.34 $\pm$0.77
FedTAD 87.08 $\pm$0.59 38.71 $\pm$1.00 89.61 $\pm$0.27 41.91 $\pm$1.14 87.15 $\pm$0.06 44.80 $\pm$0.05 93.51 $\pm$0.04 70.52 $\pm$0.53
FedSpray 86.67 $\pm$2.31 28.56 $\pm$4.74 87.40 $\pm$0.50 35.59 $\pm$0.92 86.58 $\pm$0.35 38.69 $\pm$1.36 92.01 $\pm$0.90 48.68 $\pm$1.13
FedLoG 72.74 $\pm$3.50 13.19 $\pm$3.33 79.70 $\pm$1.36 21.10 $\pm$0.85 82.25 $\pm$0.46 24.48 $\pm$1.02 OOM
DENSE 87.19 $\pm$0.44 39.18 $\pm$0.36 89.36 $\pm$0.22 41.87 $\pm$0.55 87.15 $\pm$0.07 44.92 $\pm$0.25 93.52 $\pm$0.04 70.02 $\pm$0.42
Co-Boost 87.56 $\pm$0.53 38.98 $\pm$0.82 89.52 $\pm$0.24 41.54 $\pm$0.65 87.18 $\pm$0.12 44.87 $\pm$0.30 93.54 $\pm$0.03 70.45 $\pm$0.36
O-pFGL (Ours) 88.30 $\pm$0.06 44.01 $\pm$1.93 90.24 $\pm$0.09 48.25 $\pm$1.79 90.00 $\pm$0.12 53.88 $\pm$1.15 94.21 $\pm$0.22 71.17 $\pm$1.35
Table 16: Performance of methods on more datasets under the Metis partition.
Method Computers Photo CS Physics
Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 87.50 $\pm$0.34 25.43 $\pm$0.69 90.62 $\pm$1.41 29.70 $\pm$0.16 87.39 $\pm$0.13 42.30 $\pm$0.35 93.89 $\pm$0.04 65.11 $\pm$0.20
FedAvg 87.36 $\pm$0.30 24.85 $\pm$1.69 90.76 $\pm$0.45 30.33 $\pm$1.13 87.14 $\pm$0.03 42.10 $\pm$0.05 93.65 $\pm$0.08 64.16 $\pm$0.50
FedPUB 85.20 $\pm$1.11 19.76 $\pm$1.52 89.66 $\pm$0.30 23.64 $\pm$1.85 87.30 $\pm$0.14 42.00 $\pm$0.47 93.99 $\pm$0.06 62.92 $\pm$0.96
FedGTA 74.36 $\pm$4.49 15.48 $\pm$1.66 53.68 $\pm$3.76 15.09 $\pm$1.37 87.49 $\pm$0.03 42.59 $\pm$0.11 84.00 $\pm$3.98 23.27 $\pm$0.77
FedTAD 87.64 $\pm$0.32 24.43 $\pm$0.31 90.66 $\pm$0.36 29.64 $\pm$1.44 87.15 $\pm$0.02 42.15 $\pm$0.21 93.58 $\pm$0.03 63.77 $\pm$0.17
FedSpray 86.34 $\pm$0.49 21.72 $\pm$1.82 88.77 $\pm$0.49 21.34 $\pm$0.98 81.50 $\pm$0.29 38.04 $\pm$0.26 91.17 $\pm$0.12 60.79 $\pm$0.78
FedLoG 81.68 $\pm$0.98 12.03 $\pm$1.53 85.85 $\pm$0.23 13.00 $\pm$0.22 83.79 $\pm$0.73 20.05 $\pm$1.23 OOM
DENSE 87.75 $\pm$0.29 23.99 $\pm$0.65 91.14 $\pm$0.21 31.40 $\pm$1.16 87.12 $\pm$0.08 42.04 $\pm$0.33 93.59 $\pm$0.07 64.03 $\pm$0.44
Co-Boost 87.32 $\pm$0.80 22.66 $\pm$1.74 90.97 $\pm$0.18 30.77 $\pm$0.73 87.12 $\pm$0.03 42.04 $\pm$0.08 93.57 $\pm$0.08 63.69 $\pm$0.68
O-pFGL (Ours) 88.59 $\pm$0.13 33.01 $\pm$3.66 91.79 $\pm$0.13 42.39 $\pm$0.81 89.96 $\pm$0.20 50.33 $\pm$1.01 94.51 $\pm$0.09 66.41 $\pm$1.24
A.3.7 Different GNN Architectures
To demonstrate the generality of our proposed method under different graph models, we conduct experiments with 3 other prevalent graph models, including 1) GraphSage [19], 2) GAT [45], and 3) SGC [49]. The experimental results under Louvain and Metis partitions are shown in Table 17 and Table 18, respectively. Our method shows consistent advantages over other methods.
Table 17: Performance of methods with different GNN architectures under Louvain partition.
Method GraphSage GAT SGC
Cora CiteSeer ogbn-arxiv Cora CiteSeer ogbn-arxiv Cora CiteSeer ogbn-arxiv
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 66.35 38.60 59.50 39.87 67.30 20.33 68.79 43.37 60.42 40.49 67.76 20.41 65.53 39.71 60.28 41.15 59.17 9.19
FedAvg 66.63 40.82 57.59 38.86 50.05 3.10 69.82 45.56 62.83 45.27 51.51 3.55 66.44 39.02 63.00 42.32 59.49 9.45
FedPUB 43.27 18.83 44.15 28.93 41.12 7.91 66.35 42.52 66.75 49.93 56.69 12.14 57.21 42.04 65.60 49.63 35.36 9.31
FedGTA 47.25 26.43 65.55 49.46 58.59 14.48 47.89 30.40 65.96 51.17 59.70 15.48 48.40 32.90 64.26 46.59 53.25 8.32
FedTAD 66.82 40.87 57.14 38.31 68.78 29.83 69.63 45.58 60.22 43.63 50.04 3.55 69.59 43.68 64.36 46.04 65.38 17.76
FedSpray 63.75 37.49 54.08 37.84 47.85 2.26 62.91 39.32 61.05 44.21 47.52 2.16 63.40 35.54 60.63 41.19 65.38 17.76
FedLoG 62.21 28.73 58.48 35.10 OOM 65.00 36.93 60.26 39.55 OOM N/A
DENSE 66.78 40.46 57.96 39.13 50.29 3.18 70.40 47.25 58.99 42.31 69.33 28.78 69.74 43.87 64.04 45.67 62.72 14.88
Co-Boost 65.34 39.98 54.80 35.54 68.74 29.98 70.37 47.34 59.28 42.40 69.32 28.93 68.81 44.39 64.02 45.68 62.72 14.89
O-pFGL (Ours) 72.80 54.92 69.05 54.86 69.07 32.81 75.03 57.77 71.23 57.36 70.41 34.86 73.65 54.70 69.47 54.04 70.31 35.66
Table 18: Performance of methods with different GNN architectures under Metis partition.
Method GraphSage GAT SGC
Cora CiteSeer ogbn-arxiv Cora CiteSeer ogbn-arxiv Cora CiteSeer ogbn-arxiv
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
Standalone 74.78 24.62 64.22 35.94 67.42 28.75 75.96 29.95 64.42 37.70 67.26 27.12 76.11 30.38 63.97 36.34 58.49 12.63
FedAvg 74.89 26.16 65.14 36.46 50.04 4.10 75.01 29.16 65.66 37.58 49.51 3.83 75.48 23.69 66.78 36.34 58.99 13.16
FedPUB 49.67 16.75 50.44 30.02 43.30 9.55 72.97 28.86 65.67 38.10 58.92 16.23 51.50 27.93 60.36 40.17 35.62 8.88
FedGTA 53.42 13.48 64.86 31.57 55.65 12.00 53.67 13.76 65.90 34.49 54.81 12.88 54.67 16.28 64.67 29.34 54.11 10.27
FedTAD 74.98 26.80 63.63 35.32 68.56 35.21 75.17 30.61 63.06 36.35 50.04 3.55 76.47 31.17 65.40 37.54 63.74 22.61
FedSpray 72.79 23.95 59.21 33.27 47.45 2.69 74.84 30.56 63.97 36.79 45.46 2.18 75.86 25.96 65.74 36.17 65.40 24.64
FedLoG 72.99 15.69 63.87 28.28 OOM 75.34 25.11 63.94 32.37 OOM N/A
DENSE 74.71 26.24 64.41 36.13 50.19 4.18 75.48 31.11 63.08 35.96 69.08 35.09 76.39 31.14 65.40 37.61 63.74 22.52
Co-Boost 74.56 26.69 62.27 33.17 68.53 35.08 75.44 31.09 63.32 36.27 69.14 35.09 76.42 31.15 65.42 37.62 63.74 22.52
O-pFGL (Ours) 78.48 44.59 69.55 48.17 69.20 36.57 80.68 50.02 71.91 49.03 69.98 37.83 79.96 44.48 71.58 46.87 70.26 40.10
A.3.8 Hyper-parameters Analysis
Table 19: Performance with different number of nodes per class
# nodes per class Cora CiteSeer PubMed
Acc. F1 Acc. F1 Acc. F1
1 (used in our paper) 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34
2 70.40 $\pm$2.62 54.29 $\pm$3.45 66.86 $\pm$0.56 53.63 $\pm$1.48 82.57 $\pm$0.04 66.01 $\pm$0.28
3 70.71 $\pm$1.58 52.33 $\pm$2.85 66.39 $\pm$2.73 51.61 $\pm$3.92 82.30 $\pm$0.04 65.17 $\pm$0.30
Table 20: Results on ogbn-arxiv with different number of nodes per class (percent of train set)
# nodes per class (percent of train set) ogbn-arxiv
Acc. F1
0.05% 70.51 $\pm$0.10 34.86 $\pm$0.27
0.1% 70.59 $\pm$0.15 35.28 $\pm$0.27
0.25% (used in our paper) 70.93 $\pm$0.59 36.66 $\pm$0.22
0.5% 70.98 $\pm$0.16 36.67 $\pm$0.25
1% 71.02 $\pm$0.05 36.68 $\pm$0.26
Impact of the pre-set number of nodes in synthesized global pseudo graph.
When increasing the number, the generated global pseudo-graph could have more potential to capture more information. But a large number of nodes would incur optimization difficulties and high communication costs when distributing them to clients. Thus, our principles to choose the number are (1) set the smallest number to capture enough information, (2) keep a similar label distribution with the training set. According to Section A.2, we pre-set the number of nodes to synthesize to 1 for three small graphs and pre-set the number of nodes to synthesize to $p$ percent of the total amount of nodes in corresponding class in the training set, while we keep $p$ as small as possible. We conduct experiments on 10 clients under the Louvain partition to show how the number of nodes impacts the performance. The experimental results are shown in Table 19 and Table 20. For small-scale graphs: Cora, CiteSeer, and PubMed, we pre-set the smallest number of nodes for each class (i.e., one node for each class) and find that it’s already capable. Increasing the number of nodes would not bring benefit since it incurs optimization difficulty. For larger graph: ogbn-arxiv, we pre-set the number of nodes for each class as 0.25% of the total number of corresponding class’s nodes in the train set. Fewer nodes (0.1%, 0.05%) could not fully capture the information in the aggregated class-wise statistics, and lead to a performance drop. More nodes (0.5%, 1%) are more capable but incur more computation and communication costs, and the performance improvement is marginal. Note that the number of nodes in the global pseudo-graph is always small. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150. Thus, the memory and communication costs are negligible.
Table 21: Performance with different values of $f_{th}$ and $d_{th}$
ogbn-arxiv $d_{th}=10$ $d_{th}=15$ $d_{th}=20$ $d_{th}=25$
Acc. F1 Acc. F1 Acc. F1 Acc. F1
$f_{th}=0.95$ 70.91 $\pm$0.16 36.21 $\pm$0.32 70.93 $\pm$0.59 36.66 $\pm$0.22 70.94 $\pm$0.15 36.47 $\pm$0.23 70.94 $\pm$0.08 36.37 $\pm$0.06
$f_{th}=1$ 70.82 $\pm$0.07 35.99 $\pm$0.11 70.79 $\pm$0.12 36.14 $\pm$0.51 70.72 $\pm$0.02 35.78 $\pm$0.11 70.84 $\pm$0.22 36.30 $\pm$0.61
Figure 8: Performance with different $K$ values
Table 22: Performance without $f_{th}$ and $d_{th}$
Acc. F1
Ours 76.43 $\pm$1.24 61.58 $\pm$2.16
Without $d_{th}$ (i.e., without (C1)) 75.06 $\pm$0.46 58.75 $\pm$0.82
Without $d_{th}$ and $f_{th}$ (i.e., without (C1) and (C2)) 72.51 $\pm$0.40 53.49 $\pm$2.13
Without $d_{th}$, $f_{th}$ and $topK$ (i.e., without (C1), (C2) and (C3)) 72.48 $\pm$2.45 53.48 $\pm$1.76
Impact of $d_{th}$, $f_{th}$ and $topK$ in HRE.
In our proposed Homophily-guided Reliable node Expansion strategy, we filter out the unreliable soft labels whose node degrees are less than $d_{th}$ (i.e., (C1)) and confidences are less than $f_{th}$ (i.e., (C2)). Also, we only select reliable soft labels that potentially belong to the classes with $topK$ high class homophily (i.e., (C3)). However, our method is not sensitive to the selection of the threshold $f_{th},d_{th}$, and $K$ values. We conduct experiments on ogbn-arxiv dataset under the Louvain partition with 10 clients. The experimental results under different $f_{th}$ and $d_{th}$ are shown in Table 21. We can observe that the performance remains stable under a wide range of $d_{th}$ and both $f_{th}$ values. Regarding $K$ values, each client could select a proper $K$ by statistical analysis of local label distribution, and the performance of our method is not sensitive to $K$ values. The experimental results are in Figure 8 to show the performance stability under different $K$.
To conclude, our method is not sensitive to these hyper-parameters within a proper range in HRE, and the selection would not incur complexity in deployment.
However, these criteria are still necessary because some soft labels produced by label propagation would be unreliable. We provide an experiment in Table 22 on the Cora dataset with 10 clients under the Louvain partition to demonstrate the necessity of $d_{th}$, $f_{th}$, and $topK$.
Impact of $\beta$ in Node Adaptive Distillation
In our proposed Node Adaptive Distillation, we use $\beta$ to control the scale of the value range of the $\gamma_{i}$. We provide experiments on ogbn-arxiv dataset under the Louvain partition with 10 clients to show the performance of our method under different $\beta$ values. The results are in Figure 9. We can observe that although performance would drop under too small or too large $\beta$, our method is not sensitive to $\beta$ as long as its values lie in a proper range.
Figure 9: Performance with different $\beta$ values
A.3.9 Different Adjacency Matrix Implementations
In our method, we derive the adjacency matrix $A^{\prime}$ of the generated global pseudo graph by Eq. 10. We implement the $g_{\theta}$ as a simple 3-layer MLP with input dimensions are $2\times f$ ($f$ is the node feature dimension) and output dimension is 1 (output $\in[0,1]$, represents the probability of existence of a link). Here we compare the other two different implementations of $g_{\theta}$: (1) make $A^{\prime}$ a learnable matrix and directly optimize $A^{\prime}$, (2) $A^{\prime}$ is determined by the dot product of (normalized) node features. The experimental results are shown in Table 23. The advantages of the other two designs cannot be generalized to most datasets. Our design, in contrast, offers flexibility via the learnable MLP and $\delta$ value, thus show overall performance advantages. Note that when setting $\delta=1$, our design fall back to derive a global pseudo graph where all nodes are isolated.
Table 23: Comparison of $g_{\theta}$ implementations
Cora CiteSeer PubMed ogbn-arxiv
Acc. F1 Acc. F1 Acc. F1 Acc. F1
Directly optimize $A^{\prime}$ 70.31 $\pm$4.21 54.36 $\pm$3.71 66.27 $\pm$0.80 53.12 $\pm$0.18 82.71 $\pm$0.16 66.53 $\pm$0.23 70.66 $\pm$0.06 35.54 $\pm$0.44
$A^{\prime}=\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle$ 73.01 $\pm$1.76 58.78 $\pm$3.01 69.57 $\pm$0.40 58.37 $\pm$0.08 82.68 $\pm$0.05 66.06 $\pm$0.92 70.68 $\pm$0.01 35.30 $\pm$0.54
Ours 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34 70.93 $\pm$0.59 36.66 $\pm$0.22
Table 24: Fine-grained results regarding F1-macro, Recall, and Precision metrics.
total # nodes = 339 class1 (# nodes = 50) class2 (# nodes = 14)
F1-macro Recall Precision F1-macro Recall Precision
Standalone 51.28 50.00 52.63 0 0 0
FedAvg 45.71 40.00 53.33 0 0 0
FedPUB 55.81 60.00 52.17 0 0 0
FedGTA 50.00 55.00 45.83 29.63 57.14 20.00
FedTAD 42.42 35.00 53.85 0 0 0
FedSpray 50.00 56.25 45.00 0 0 0
FedLoG 47.06 40.00 57.14 0 0 0
DENSE 48.65 45.00 52.94 0 0 0
Co-Boost 42.42 35.00 53.85 0 0 0
O-pFGL (Ours) 63.83 75.00 55.56 50.00 42.86 60.00
A.3.10 Fine-grained Experimental Results
In Figure 1, we’ve separately compared the F1-macro of our and baseline methods on minor classes (class 1 and class 2). The class-specific F1 and F1-macro together support that our method not only improves the performance of major classes, but can also effectively generalize to the minority. We select one client among 10 clients on the CiteSeer dataset under the Louvain partition to provide more comparisons on minor classes regarding F1-macro, Recall, and Precision metrics in Table 24.
A.4 Complexity Analysis
We provide the complexity analysis of methods from three perspectives: communication, computation, and memory, with a 2-layer GCN model as an example.
Table 25: Communication complexity of methods.
Communication Upload Download
Standalone N/A N/A
FedAvg $O(fd+dC)$ $O(fd+dC)$
FedPUB $O(fd+dC)$ $O(fd+dC)$
FedGTA $O(fd+dC+kKC)$ $O(fd+dC)$
FedTAD $O(fd+dC)$ $O(fd+dC)$
FedSpray $O(fd^{\prime}+d^{\prime}C)$ $O(fd^{\prime}+d^{\prime}C)$
FedLoG $O(fd+dC+n^{\prime 2}f+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t})$ $O(fd+dC+\text{Clf}_{h}+\text{Clf}_{t}+n^{\prime 2}f)$
DENSE $O(fd+dC)$ $O(fd+dC)$
Co-Boost $O(fd+dC)$ $O(fd+dC)$
O-pFGL (Ours) $O(Chf)$ $O(n^{\prime}f+n^{\prime 2})$
Communication Analysis.
We provide the communication analysis of the methods in Table 25. FedAvg, FedPUB, FedTAD, DENSE, and Co-Boost all need to upload and download the whole model parameters ($O(fd+dC)$, where $d$ is the hidden dimension). FedGTA needs to upload additional mixed moments of neighboring features ($O(kKC)$, where $K$ is the order of the moment and $k$ is the propagation steps of Label Propagation) for personalized aggregation. FedSpray needs to upload and download the global feature-structure encoder and structure proxy ($O(fd^{\prime}+d^{\prime}C)$, where $d^{\prime}$ is the output dimension of the encoder and dimension of the structure proxy). FedLoG needs to upload the pre-trained prompt generator (for simplicity, we use $O(\text{PG})$ to denote the size of the prompt generator), head-branch classifier and tail-branch classifier (we use $O(\text{Clf}_{h})$ and $O(\text{Clf}_{t})$ to denote the size of the two branch classifiers), synthetic graph ($O(n^{\prime 2}f)$, where $n^{\prime}$ is the pre-set number of nodes), along with the model, to the server. Each client in FedLoG needs to download the aggregated head-branch and tail-branch classifier and the model. The PG, Clfh, and Clft are implemented as multi-layer perceptrons with relatively large hidden dimensions. Thus, FedLoG would incur high upload and download communication costs. In contrast, our method only needs to upload the class-wise statistics ($O(chf)$) and download the generated global pseudo-graph ($O(n^{\prime}f+n^{\prime 2})$). $C$ is the number of classes. $h$ is the propagation step in Eq. 5 and $h\leq 2$ in our experiments. Although in some rare cases where $C$ is extremely large, the upload costs still would not become a huge burden compared with the model size. $n^{\prime}$ is the number of nodes in the global pseudo-graph, which is usually a pre-set small value, as indicated in Sec A.2. The communication costs of our method are independent of the model size. The larger the model, the more communication-efficient our method is compared to other methods.
Table 26: Computation complexity of methods.
Computation Client Server
Standalone $O(m(f+d)+nd(f+C))$ N/A
FedAvg $O(m(f+d)+nd(f+C))$ $O(N(fd+dC))$
FedPUB $O(m(f+d)+nd(f+C))$ $O(N^{2}(fd+dC)+Kn^{\prime 2})$
FedGTA $O(m(f+d+kC)+nd(f+C))$ $O(N^{2}(fd+dC+kKC))$
FedTAD $O(pn^{2}+m(f+d)+nd(f+C))$ $O(N(fd+dC)+n^{\prime 2}f+N(m^{\prime}(f+d)+n^{\prime}(d(f+C)))+L^{\prime}n^{\prime}d^{\prime 2})$
FedSpray $O(m(f+d)+nd(f+C)+nfd^{\prime}+nd^{\prime}C)$ $O(Nfd^{\prime})$
FedLoG $O(m(f+d)+nd(f+C)+(n^{\prime}+n)d^{\prime 2})$ $O(N(fd+dC+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t}+n^{\prime 2}f))$
DENSE $O(m(f+d)+nd(f+C))$ $O(N(fd+dC)+n^{\prime 2}f+N(m^{\prime}(f+d)+n^{\prime}(d(f+C)))+L^{\prime}n^{\prime}d^{\prime 2})$
Co-Boost $O(m(f+d)+nd(f+C))$ $O(N(fd+dC)+n^{\prime 2}f+N(m^{\prime}(f+d)+n^{\prime}(d(f+C)))+L^{\prime}n^{\prime}d^{\prime 2})$
O-pFGL (Ours) $O(m(f+d)+nd(f+C))$ $O(NChf+m^{\prime}f+Chf+n^{\prime 2}f)$
Computation Analysis.
We provide computation analysis of methods in Table 26. We slightly abuse the notation by redefining the number of clients as $N$, the number of edges in local graph data as $m$, and the number of nodes in local graph data as $n$. $K^{\prime}$ is the number of proxy graphs generated in FedPUB, and $n^{\prime}$ is the number of nodes in each proxy graph. $p$ is the diffusion step in FedTAD, and $O(pn^{2})$ is the complexity to calculate the topology embedding, which could be highly computationally expensive. For FedTAD, DENSE, and Co-Boost, $O(n^{\prime 2}f)$ is the complexity of constructing the generated graph. $L^{\prime}$ is the number of layers of the generator, $f^{\prime}$ is the dimension of the latent features in the generator, and $n^{\prime}$ is the number of generated nodes. $O(L^{\prime}n^{\prime}d^{\prime 2})$ is the complexity of forwarding the generator in the server. For FedLoG, the additional computation costs come from the prompt generator pretraining and graph synthesis. Although the prompt generator and classifiers are multilayer perceptrons, the pre-training process involves bilevel optimization, which would incur high computation costs in practice. For our method, $O(NChf)$ is the complexity of aggregating the statistics. $m^{\prime}$ is the estimated number of edges in the generated global pseudo-graph, usually a small value. $O(m^{\prime}f+Chf)$ is the complexity of calculating alignment loss. $O(n^{\prime 2}f)$ is the complexity of calculating the smoothness loss. The computation complexity on clients in our method is kept as least as possible, the same as the vanilla FedAvg and Standalone.
Table 27: Memory complexity of methods.
Memory Client Server
Standalone $O(fd+dC)$ N/A
FedAvg $O(fd+dC)$ $O(N(fd+dC))$
FedPUB $O(fd+dC)$ $O(N(fd+dC)+K^{\prime}n^{\prime}f)$
FedGTA $O(fd+dC+kKc)$ $O(N(fd+dC+kKc))$
FedTAD $O(fd+dC+pn^{2})$ $O(N(fd+dC)+n^{\prime}f+n^{\prime 2})$
FedSpray $O(fd+dC+fd^{\prime}+d^{\prime}C+nd^{\prime})$ $O(Nnd^{\prime}+f^{\prime}d^{\prime}+d^{\prime}C)$
FedLoG $O(fd+dC+n^{\prime}f+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t}+n\overline{d}^{k}f)$ $O(N(fd+dC+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t}+n^{\prime 2}f))$
DENSE $O(fd+dC)$ $O(N(fd+dC)+n^{\prime}f+n^{\prime 2})$
Co-Boost $O(fd+dC)$ $O(N(fd+dC)+n^{\prime}f+n^{\prime 2})$
O-pFGL (Ours) $O(fd+dC)$ $O(Nchf+n^{\prime}f+n^{\prime 2})$
Memory Analysis.
We provide memory analysis of methods in Table 27. Standalone, FedAvg, FedPUB, DENSE, and Co-Boost only need to store the model parameters ($O(fd+dC)$). FedTAD needs to additionally store the topology embedding calculated by graph diffusion ($O(pn^{2})$), thus it faces OOM on larger graphs. FedSpray needs to store additional encoders and structure proxy. FedLoG needs to store the additional prompt generator, two branch classifiers, and generated graphs. Notably, in prompt generator pretraining, FedLoG needs additionally obtain the k-hop subgraph of each node in the local graph, which would incur huge memory ($O(n\overline{d}^{k}f)$, where $\overline{d}$ is the average degree and $k$ is the pre-set number of hops). This is the key factor leading to the out-of-memory issue in dense graphs or large-scale graphs. For FedTAD, DENSE, Co-Boost, and our method, the server needs to additionally store the generated graph, which costs approximately $O(n^{\prime}f+n^{\prime 2})$ memory. Note that $n^{\prime}$ is always a small number. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150 ($n^{\prime}<150$). Thus, it would not incur much memory overhead on the server.
A.5 An Equivalent Server-Efficient Variant Method
An equivalent variant of our method is to move the global pseudo-graph generation process to the clients. Specifically, the server only securely aggregates the uploaded statistics and then distributes $\{N^{c},\boldsymbol{\mu}^{c},\boldsymbol{s^{2}}^{c}\}$ to clients (with Homomorphic Encryption [1]). Then the clients generate the global pseudo-graph with aggregated statistics locally. It still supports model heterogeneity and Secure Aggregation protocols. This variant enables extreme efficiency on the server since the server only needs to perform a series of weighted average operations, moving the computation overhead of global pseudo-graph generation to the client-side.
A.6 Related Works
A.6.1 Federated Graph Learning
With the rapid development of federated learning methods, recent works introduce federated graph learning to collaboratively train graph models [31] in a privacy-preserving manner and apply it to many applications [48, 61, 53, 44, 7]. From the graph level, each client possesses multiple completely disjoint graphs (e.g., molecular graphs). Recent works [52, 43] mainly focus on the intrinsic heterogeneity among graphs from different clients. From the subgraph level, the graph possessed by each client can be regarded as a part of a larger global graph. To cope with heterogeneous graph data, GraphFL [46] adopts meta-learning [12] for better generalization. FedGL [5] uploads the node embedding and prediction for global supervision but faces a heavy communication burden and potential privacy concerns. FGSSL [24] augments the local graph to mitigate the heterogeneity. To enhance the model utility on each client, personalized federated graph learning methods are proposed. FedPUB [2] generates random graphs [22] to measure the similarity in model aggregation and conducts adaptive weight masks for better personalization. FedGTA [33] proposes topology-aware personalized optimization. AdaFGL [32] studies the structure non-IID problem. FedTAD [69] utilized ensemble local models to perform data-free distillation on the server. To complete missing connections between graphs of clients, FedSage [63] and FedDEP [64] additionally train a neighborhood generator. FedGCN [57] additionally uploads and downloads the encrypted neighbor features to supplement the features of the neighborhood. To better represent the minority in local graph data, FedSpray [15] learns local class-wise structure proxies to mitigate biased neighboring information. But it needs numerous communication rounds for optimization. FedLoG [30] synthesizes node features in each client, and the server aggregates these node features to generate global synthetic data. Each client then generalizes its local training via the synthetic graph.
A.6.2 One-shot Federated Learning
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into 3 categories. (1) Ensemble-based: The original OFL study [18] ensembles local models and conducts knowledge distillation with public data. DENSE [62] employs model inversion [58] from the ensemble model to generate images for distillation. FedOV [10] introduces placeholders in the model prediction layer. IntactOFL [60] trains a MoE [27] network by the generated images. Co-Boost [9] further optimizes the generated images and ensemble weights iteratively. (2) Distillation-based: DOSFL [66] and FedD3 [42] conduct dataset distillation locally and upload distilled data for server-side training. (3) Generative-based: FedCVAE [21] trains VAEs for each client to generate similar images, addressing data heterogeneity. FedDISC [54] and FedDEO [55] leverage the pre-trained Stable Diffusion [38] to generate images and mitigate data heterogeneity. However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning.
A.6.3 Critical Analysis of Current Related Work
Existing FGL methods require numerous communication rounds.
Existing FGL methods typically follow the conventional paradigm in which the server adaptively aggregates client models. Under practical non-IID scenarios, numerous studies have shown that this paradigm can lead to issues such as weight divergence [65], client drift [29], and biases and conflicts among client models [70]. These challenges necessitate extensive communication rounds to achieve acceptable performance.
Specialized FGL methods (e.g., FedPUB, FedTAD, FedGTA) necessitate unbiased client models with strong predictive abilities for various tasks, including proxy embedding computation, knowledge distillation, and label propagation. These methods require significant communication rounds to resolve these challenges. FedSpray additionally optimizes the structure proxy, which demands increased communication rounds for optimization. FedLoG needs to optimize the global synthetic data, which also needs iterative communications.
Existing OFL methods are not designed for graph data.
Existing OFL methods are mostly designed for image data. When dealing with graph data, these methods are incompatible or impractical. To be specific, ensemble-based methods such as DENSE train a generator to generate pseudo images via model inversion. However, it’s not perfectly compatible with graph data generation since it can hardly capture the fine-grained structural information of nodes. Distillation-based methods generate distilled data on clients, and the server assembles the distilled data. However, the nodes in a graph are not independent. It’s tricky to assemble generated graphs from clients on the server side. Generative-based methods use VAE or pretrained Stable Diffusion, which cannot generate graph data. Thus, there are not OFL methods that fit perfectly for graph data.
Most FGL and OFL methods are not compatible with Secure Aggregation.
Secure Aggregation protocols are widely used in federated learning, which enable the server to compute the sum of large, user-held data vectors securely, without accessing the individual client contributions. Thus, the client’s uploaded models or other information could be protected from third-party and the curious server. The application of Secure Aggregation is based on the fact that only weighted average operations are conducted in vanilla federated learning methods like FedAvg. However, the recent federated graph learning methods have become complicated. The server needs to conduct other operations, including ensemble (e.g., FedTAD, DENSE, and Co-Boost), data-free distillation (e.g., FedTAD, DENSE, and Co-Boost), functional embedding calculation (e.g., FedPUB), complex similarity calculation (e.g., FedGTA), and others (e.g., FedLoG). The server must access every client uploading to conduct these operations. Thus, they are not compatible with Secure Aggregation and third-party, and the curious server could easily access each client’s uploaded information.
A.7 Baselines.
We provide a detailed description of the baseline methods in the experiments.
Standalone is a non-federated learning method where each client trains its model only use its local graph data, without communication.
FedAvg is the fundamental federated learning method where each client trains its local model and upload to the server. The server weighted average the client models’ parameters to obtain the global model and then distributes the global model to clients. In our experiments, we limit the upload-download communication round to one round. For better performance, we perform additional finetuning on local graph data after downloading the global model.
FedPUB is a personalized federated graph learning method. FedPUB computes similarities between local models via functional embeddings derived from random graphs generated in the server. The similarities are used for weighted average in personalized aggregation. Additionally, FedPUB employs a personalized mask for each client model to selectively upload parameters. In our experiments, we limit the upload-download communication round to one round.
FedGTA is a personalized federated graph learning method. FedGTA compute and upload the local smoothing confidence and mixed moments of neighbor features to the server to facilitate the weighted average in personalized aggregation. In our experiments, we limit the upload-download communication round to one round.
FedTAD computes the class-wise knowledge reliability scores on each client. The scores, along with local models, are uploaded to the server for data-free knowledge distillation from the ensemble of local models to the global model or personalized model. In our experiments, we limit the upload-download communication round to one round. For better performance, we perform additional finetuning on local graph data after downloading the model.
FedSpray focuses on the minor classes. It optimizes the structure proxy along with the local model on each client. FedSpray uploads and aggregate the structure proxy, which is used in local model training to prevent overfitting and ensure the generalization to the minor classes. In our experiments, we limit the upload-download communication round to one round.
FedLoG focuses on the local generalization of model training. Each client pretrains a prompt generator and synthesizes nodes. The server aggregates the model and synthesized nodes and generates global synthetic data. Then, each client trains its model on local graph data and global synthetic data to derive a generalized model. In our experiments, we limit the upload-download communication round to one round.
DENSE is a one-shot federated learning method. Each client uploads its trained local model to the server. The server ensembles the client models to generate pseudo data via model inversion, which are used to distill knowledge from client models to the global model or personalized models. For better performance, we perform additional finetuning on local graph data after downloading the data. Note that it’s designed for the image classification task. In our node classification task on graphs, the generators in model inversion on the server-side are trained to generate node features, and the topology structure is constructed using the $K$-Nearest Neighbors strategy as outlined in [69].
Co-Boost is a one-shot federated learning method. It advances the DENSE. The server ensembles the client models with learnable weights. The ensemble weights and generated pseudo data are optimized alternately. For better performance, we perform additional finetuning on local graph data after downloading the data. Note that it’s designed for the image classification task. In our node classification task on graphs, the generators in model inversion on the server-side are trained to generate node features, and the topology structure is constructed using the $K$-Nearest Neighbors strategy as outlined in [69].
A.8 Our method’s Advantages in One Communication Round
Most existing methods meet issues like weight divergence, client drift, and discrepancies and conflicts between client models raised from non-IID scenarios and fewer communication rounds. However, our method is robust to non-IID and can even achieve optimal performance in one round on average compared with multi-round federated learning methods. In our method, we do not aggregate the model parameters (which is a major source of information loss in other methods). Rather, the aggregation process in our method is lossless and unbiased regardless of the data distributions. Thus, we do not need multi-round communications to tackle the issues met by the existing methods above. Also, our method enhances the local generalization of the client models by considering both local and global information, which further boosts our method’s performance.
A.9 Experimental Results on a Large Number of Clients
We conduct experiments on a large number of clients to evaluate the robustness of our method. We simulate 10, 100, and 1000 clients on ogbn-products with Metis partition. The F1-macro metrics of methods are shown in Figure 10. Our method consistently outperforms other methods on a large number of clients.
Figure 10: Performance of method with a large number of clients on ogbn-products dataset under the Metis partition.
(a) Cora
(b) CiteSeer
(c) PubMed
Figure 11: Performance of method on the subsequent link prediction task on Cora, CiteSeer, PubMed datasets. The x-axis is set to a logarithmic scale.
A.10 Experimental Results on Link Prediction Task
To evaluate the effectiveness of the methods of node representation learning, we perform the link prediction task. We first perform the node classification task to pre-train a GCN as the feature encoder with a classification head to generate distinguishable node representations, and then fine-tune the feature encoder for link prediction. Note that FedLoG is not included since its inference needs the participation of synthetic head and tail branch nodes, which cannot apply to the link prediction task directly. The experimental results (take AUC as our metrics) are shown in Figure 11, which shows that our method learns better node representation and achieves the best on the subsequent link prediction task while having less communication cost.
A.11 Limitations
Our method is based on the assumption that the clients and the server are honest and benign in the federated learning process, which is a common assumption in this research area. We leave the discussion about border scenarios with fewer assumptions for future work.